### v1

# Nonuniqueness of capped black holes: large and small bubbles

Abstract
We present a new non-BPS solution describing an asymptotically flat, stationary, bi-axisymmetric capped black hole in the bosonic sector of five-dimensional minimal supergravity. This solution describes a spherical black hole, while the exterior region of the horizon exhibits a non-trivial topology of $[{\mathbb{R}}^{4}\#{\mathbb{C}}{\mathbb{P}}^{2}]\setminus{\mathbb{B}}^{4}$ on a timeslice. This solution extends our previously constructed three-parameter solution to a more general four-parameter solution. To derive this solution, we utilize a combination of the Ehlers and Harrison transformations and then impose appropriate boundary conditions on the solution’s parameters. It can be shown that the resultant solution is free from curvature, conical, Dirac-Misner string and orbifold singularities, as well as closed timelike curves on and outside the horizon. Characterized by four independent conserved charges—mass, two angular momenta, and electric charge—this solution reveals two distinct branches: a small bubble branch and a large bubble branch, distinguished by non-conserved local quantities such as magnetic flux or magnetic potential. This shows the non-uniqueness for spherical black holes, even among capped black holes. For equivalent sets of conserved charges, we find that the large/small bubble branch can have larger/smaller entropy than the Cvetič-Youm black hole.
preprint: TTI-MATHPHYS-32
I Introduction
The topological censorship theorem imposes significant constraints on the allowed topology of the domain of outer communication (DOC) Friedman:1993ty . In four dimensions, only a trivial topology represented by ${\rm DOC}\cap\Sigma=\mathbb{R}^{3}\setminus\mathbb{B}^{3}$ ($\Sigma$: timeslice) is permissible. However, in higher dimensions, non-trivial topologies that may have homology groups of higher rank are allowed. From this theorem, Ref. Hollands:2010qy demonstrated that in a five-dimensional stationary and axisymmetric spacetime, ${\rm DOC}\cap\Sigma$ can admit the non-trivial topology of $[\mathbb{R}^{4}\#n(\pm\mathbb{C}{\mathbb{P}}^{2})\#m(S^{2}\times S^{2})]\setminus\mathbb{B}^{4}$. Despite these possibilities, in higher dimensions for static asymptotically flat spacetimes, no black hole solution with such nontrivial DOC exists due to the uniqueness theorems Gibbons:2002bh ; Gibbons:2002av . These theorems assert that the higher-dimensional Schwarzschild and Reissner-Nordström solutions Tangherlini:1963bw are the only vacuum and charged black hole solutions, respectively. The uniqueness theorem for charged rotating black holes in the bosonic sector of five-dimensional minimal supergravity Tomizawa:2009ua states that under the assumptions of two commuting axial isometries and spherical topology of horizon cross-sections, an asymptotically flat, stationary charged rotating black hole with a non-extremal horizon is uniquely characterized by its mass, charge, and two independent angular momenta and is therefore described by the five-dimensional Cvetič-Youm solution Cvetic:1996xz . However, unlike the uniqueness theorem in four dimensions, this theorem assumes the additional condition that the intersection of a black hole’s exterior region with the timeslice $\Sigma$ has a trivial topology, $\mathbb{R}^{4}\setminus\mathbb{B}^{4}$. Thus, it is still possible that such a solution with DOC of nontrivial topologies exists, and if it exists, it cannot be static but must belong to a class of stationary solutions.
Kunduri and Lucietti have constructed a BPS (Bogomol’nyi-Prasad-Sommerfield) black hole solution within the bosonic sector of five-dimensional minimal supergravity, which features the DOC of the topology $[{\mathbb{R}}^{4}\#(S^{2}\times S^{2})]\setminus{\mathbb{B}}^{4}$ Kunduri:2014iga . In our previous work, we developed the first non-BPS exact solution representing an asymptotically flat, stationary spherical black hole with a nontrivial topology of $[\mathbb{R}^{4}\#\mathbb{CP}^{2}]\setminus\mathbb{B}^{4}$ in five-dimensional minimal supergravity. This solution was derived using a methodology similar to that employed for the charged dipole black ring previously constructed in our work Suzuki:2024coe . Initially, we utilized the inverse scattering method (ISM) to construct a vacuum black hole possessing a Dirac-Misner string singularity. Following this, we applied the electric Harrison transformation to this vacuum solution, resulting in a charged rotating solution characterized by an $S^{3}$ horizon topology, yet still retaining the Dirac-Misner string singularity. To ensure the regularity of the solution, we finely tuned the solution’s parameters to eliminate the Dirac-Misner string singularity. The final solution has independent three parameters and exhibits regularity, with the absence of curvature, conical, Dirac-Misner string, and orbifold singularities, both inside and outside the horizon, as well as the absence of closed timelike curves (CTCs). It describes a charged rotating black hole capped by a disk-shaped bubble, hence we called it “capped black hole”.
In this paper, using the Ehlers-Harrison transformations, previously applied to the charged rating black ring with all independent conserved charges Suzuki:2024vzq , we wish to generalize the three-parameter family of the capped black hole solution, derived in our previous works Suzuki:2023nqf ; Suzuki:2024phv , to a four-parameter family. The Ehlers and Harrison transformations add an angular momentum and an electric charge to a five-dimensional vacuum solution, respectively, preserving asymptotic flatness. In our previous work, we developed the construction of a vacuum black ring solution possessing a Dirac-Misner string singularity through the Ehlers transformation, demonstrating that a doubly rotating black ring (Pomeransky-Sen’kov solution) can be derived once the Dirac-Misner string singularity is removed. Subsequently, in Ref. Suzuki:2024vzq , we applied the Harrison transformation to this vacuum solution, generating a corresponding charged solution within five-dimensional minimal supergravity. By setting appropriate boundary conditions on the parameters, we successfully eliminated all singularities—including those related to curvature, conical features, and Dirac-Misner strings—across the rotational axes and the horizon. Using this methodology, we can construct a capped black hole, distinguishing it by replacing the $S^{1}\times S^{2}$ topology with $S^{3}$. We demonstrate that the new solution manifests discrete non-uniqueness in terms of conserved charges, presenting two distinct branches that share the same mass, angular momenta, and electric charges as the Cvetič-Youm black hole. These branches are distinguished by the size of their bubbles adjacent to the horizon and are accordingly termed small bubble and large bubble branches. Our analysis indicates that within one parameter region, the large bubble branch exhibits greater entropy than both the small bubble branch and the Cvetič-Youm black hole, making it thermodynamically more favored. Conversely, in another parameter region, the Cvetič-Youm black hole displays higher entropy than both the large and small bubble branches.
The rest of the paper is organized as follows. In section II, we briefly review the derivation of the solution for the metric and gauge potential, and present their expression in the C-metric coordinates. In section III, we clarify the boundary condition for the absence of curvature, conical, orbifold, Dirac-Misner string singularities and closed timelike curves. In section IV, we investigate the phase of the physical solution. Finally, we summarize the results in section V.
II Solution generation
II.1 Setup
We consider asymptotically flat, stationary, and bi-axisymmetric solutions in the bosonic sector of five-dimensional minimal ungauged supergravity (Einstein-Maxwell-Chern-Simons theory), whose the action is expressed as follows
$S=\frac{1}{16\pi G_{5}}\left[\int d^{5}x\sqrt{-g}\left(R-\frac{1}{4}F^{2}\right)-\frac{1}{3\sqrt{3}}\int F\wedge F\wedge A\right]\,,$ (1)
where $F=dA$. The field equation comprises of the Einstein equation
$R_{\mu\nu}-\frac{1}{2}Rg_{\mu\nu}=\frac{1}{2}\left(F_{\mu\lambda}F_{\nu}^{~{}\lambda}-\frac{1}{4}g_{\mu\nu}F_{\rho\sigma}F^{\rho\sigma}\right)\,,$ (2)
and the Maxwell equation with a Chern-Simons term
$d\star F+\frac{1}{\sqrt{3}}F\wedge F=0\,.$ (3)
Assuming that there is one timelike Killing vector $\xi_{0}=\partial/\partial t$ and one spacelike axial Killing vector $\xi_{1}=\partial/\partial\psi$, this theory simplifies to the $G_{2(2)}/SL(2,{\mathbb{R}})\times SL(2,{\mathbb{R}})$ non-linear sigma models coupled with three-dimensional gravity Mizoguchi:1998wv ; Mizoguchi:1999fu . Following the discussion in Ref. Tomizawa:2009ua , from Eqs. (2) and (3), one can define the gravitational potential $\lambda_{ab}$, electric potential $\psi_{a}$, magnetic potential $\mu$ and twist potentials $\omega_{a}$ by
$\lambda_{ab}$ $=$ $(\xi_{a}|\xi_{b}),$ (4)
$d\psi_{a}$ $=$ $-i_{\xi_{a}}F,$ (5)
$d\mu$ $=$ $\frac{1}{\sqrt{3}}\star(\xi_{0}\wedge\xi_{1}\wedge F)-\epsilon^{ab}\psi_{a}d\psi_{b}\,,$ (6)
$d\omega_{a}$ $=$ $\star(\xi_{0}\wedge\xi_{1}\wedge d\xi_{a})+\psi_{a}(3d\mu+\epsilon^{bc}\psi_{b}d\psi_{c})\,,$ (7)
where $\epsilon^{01}=-\epsilon^{10}=1$. Then, the action (1) reduces to the nonlinear sigma model for the eight scalar functions $\{\lambda_{ab},\omega_{a},\psi_{a},\mu\}$ coupled with three-dimensional gravity.
Furthermore, if we assume the existence of a third spacelike axial Killing vector $\xi_{2}=\partial/\partial\phi$ commuting with the other two, the metric can be expressed in the Weyl-Papapetrou form
$ds^{2}$ $=$ $\lambda_{ab}(dx^{a}+a^{a}{}_{\phi}{}d\phi)(dx^{b}+a{}^{b}{}_{\phi}d\phi)+\tau^{-1}\rho^{2}d\phi^{2}+\tau^{-1}e^{2\sigma}(d\rho^{2}+dz^{2})\,,$ (8)
and the gauge potential is written,
$A=\sqrt{3}\psi_{a}dx^{a}+A_{\phi}d\phi\,,$ (9)
where the coordinates $x^{a}=(t,\psi)$ ($a=0,1$) denote the Killing coordinates, and thus all functions $\lambda_{ab}$, $\tau:=-{\rm det}(\lambda_{ab})$, $a^{a}$, $\sigma$, and $(\psi_{a},A_{\phi})$ are independent of $\phi$ and $x^{a}$. Note that the coordinates $(\rho,z)$ that span a two-dimensional base space, $\Sigma=\{(\rho,z)|\rho\geq 0,\ -\infty<z<\infty\}$, are globally well-defined, harmonic, and mutually conjugate on $\Sigma$. The metric functions $a^{a}{}_{\phi}\ (a=0,1)$ and the component $A_{\phi}$ of the gauge potential are determined by the eight scalar functions $\{\lambda_{ab},\omega_{a},\psi_{a},\mu\}$ from Eqs. (6) and (7), and the function $\sigma$ is determined from these scalar functions up to a constant factor. Therefore, the theory reduces to the nonlinear sigma model of $\{\lambda_{ab},\omega_{a},\psi_{a},\mu\}$ decoupled from the gravity. Because the theory is invariant under the $G_{2(2)}$-transformation, one can produce a new solution from a known solution by applying a certain transformation in the $G_{2(2)}$-group. Here, we focus on two different $G_{2(2)}$-transformations.
One is a transformation in the gravitational sector that reduces to $SL(3,{\mathbb{R}})/SL(2,{\mathbb{R}})$ non-linear sigma model for the vacuum solution $\Phi^{A}=\{\lambda_{ab},\omega_{a},\psi_{a}=0,\mu=0\}$, which is invariant under the $SL(3,{\mathbb{R}})$ transformation Maison:1979kx . Giusto and Saxena found that a $SL(2,{\mathbb{R}})$ subgroup in $SL(3,{\mathbb{R}})$, or the Ehlers transformation, can be utilized to add angular momenta to five-dimensional asymptotically-flat, vacuum solutions Giusto:2007fx . With this transformation, one can obtain the new metric describing the rotating solution to the vacuum Einstein equations with the same Killing isometries. For example, in Ref. Giusto:2007fx , applying this transformation to the five-dimensional static black hole solution (the five-dimensional Schwarzschild-Tangherlini solution Tangherlini:1963bw ) one can reproduce the five-dimensional rotating black hole solution (the five-dimensional Myers-Perry solution Myers:1986un ). However, when applied to vacuum black rings, such as the static black ring and the singly rotating black ring, a Dirac-Misner string singularity inevitably appears on the disk inside the ring, similar to the Harrison transformation studied in Ref. Bouchareb:2007ax . This issue is solved by using the seed with a Dirac-Misner string singularity. In our prior study Suzuki:2024eoz , applying the Ehlers transformation to the vacuum solution with a Dirac-Misner string singularity constructed in Ref. Chen:2008fa , we can successfully reproduce the known doubly-rotating vacuum black ring solution (Pomeransky-Senkov solution) Pomeransky:2006bd .
The other is the so-called Harrison transformation formulated by Bouchareb et. al. in Ref. Bouchareb:2007ax . This transforms a five-dimensional vacuum solution $\Phi^{A}=\{\lambda_{ab},\omega_{a},\psi_{a}=0,\mu=0\}$ into a charged solution $\Phi^{\prime A}=\{\lambda^{\prime}_{ab},\omega^{\prime}_{a},\psi^{\prime}_{a},\mu^{\prime}\}$ in the five-dimensional minimal supergravity while preserving asymptotic flatness. In particular, this transformation adds the electric and dipole charges to a vacuum black ring solution while keeping the asymptotic flatness and Killing isometries. However, as mentioned in Ref. Bouchareb:2007ax , when one performs the Harrison transformation for the regular vacuum black ring such as the Pomeransky-Sen’kov solution, a Dirac-Misner string singularity inevitably appears on the disk inside the ring, though the transformation can generate the regular Cvetič-Youm charged black hole for the vacuum black hole such as the Myers-Perry solution. Instead, performing the Harrison transformation to the vacuum solution with a Dirac-Misner string singularity, one can obtain regular charged black holes such as the capped black hole Suzuki:2023nqf ; Suzuki:2024phv and charged black rings Suzuki:2024coe ; Suzuki:2024vzq .
II.2 Charged solution
We aim to obtain a generalized version of the capped black hole obtained in Ref. Suzuki:2023nqf ; Suzuki:2024phv . To this end, we prepare the charged solution by successively applying the Ehlers and Harrison transformation to the vacuum seed solution produced by the inverse scattering method in Ref. Chen:2008fa . This charged solution was first constructed in our previous work Suzuki:2024vzq , to derive the charged black ring solution. We show that the same solution also contains the capped black hole under the different boundary condition. Below, we start with revisiting the charged solution obtained in Ref. Suzuki:2024vzq , whose metric and gauge potential are given by
$ds^{2}=-\frac{H(y,x)}{D^{2}H(x,y)}(dt+\Omega^{\prime}_{\psi}(x,y)d\psi+\Omega^{\prime}_{\phi}(x,y)d\phi)^{2}+\frac{D}{H(y,x)}\left[F(y,x)d\psi^{2}-2J(x,y)d\psi d\phi-F(x,y)d\phi^{2}\right]$
$\quad+\frac{\ell^{2}DH(x,y)}{4(1-\gamma)^{3}(1-\nu^{2})(1-a^{2})\Delta(x-y)^{2}}\left(\frac{dx^{2}}{G(x)}-\frac{dy^{2}}{G(y)}\right),$ (10)
and
$A=\frac{\sqrt{3}cs}{DH(x,y)}\left[(H(x,y)-H(y,x))dt\right.$
$\hskip 56.9055pt\left.-(cH(y,x)\Omega_{\psi}(x,y)-sH(x,y)\Omega_{\phi}(y,x))d\psi-(cH(y,x)\Omega_{\phi}(x,y)-sH(x,y)\Omega_{\psi}(y,x))d\phi\right],$ (11)
with
$D$ $:=\frac{c^{2}H(x,y)-s^{2}H(y,x)}{H(x,y)},$ (12)
$\Omega_{\psi}^{\prime}(x,y)$ $:=c^{3}\Omega_{\psi}(x,y)-s^{3}\Omega_{\phi}(y,x),$ (13)
$\Omega_{\phi}^{\prime}(x,y)$ $:=c^{3}\Omega_{\phi}(x,y)-s^{3}\Omega_{\psi}(y,x),$ (14)
where the parameter $\alpha$ induced by the Harrison transformation appears as $(c,s):=(\cosh\alpha,\sinh\alpha)$. The metric functions become
$G(u)$ $=(1-u^{2})(1+\nu u),$ (15)
$H(x,y)$ $=(1+y)^{2}\biggr{[}\nu d_{1}f_{7}^{2}(1-\nu)\left(1-x^{2}\right)\left(2(\gamma-1)\nu+c_{3}\right)+2c_{3}c_{1}^{2}f_{6}^{2}(\gamma-\nu)(\nu x+1)^{2}$
$\quad-(\gamma-\nu)(1-\nu)\left(c_{1}^{2}c_{3}f_{6}^{2}-2(1-\gamma)\nu\left(c_{1}-b(1-\nu)\right){}^{2}f_{8}^{2}\right)(1-x)(1+x\nu)\biggr{]}$
$-(1+y)\biggr{[}d_{5}g_{6}(1-x)^{2}+d_{6}g_{7}\left(1-x^{2}\right)+2(1-\nu)^{-1}d_{6}g_{8}(1+x)(1+x\nu)\biggr{]}$
$+4\left(1-a^{2}\right)(1-x)(1-\gamma)^{3}(1-\nu)^{4}\Delta-4(1+x)(1-\gamma)(1-\nu)^{2}(1+\nu)d_{1}f_{2}^{2}+2\left(1-x^{2}\right)(1-\gamma)(1-\nu)^{2}c_{2}^{2}f_{5}^{2},$ (16)
$F(x,y)$ $=\frac{\ell^{2}\tilde{v}_{0}^{2}(1-\gamma)}{(\gamma^{2}-\nu^{2})(x-y)^{2}}\biggr{[}4(1+y\nu)G(x)\left\{\left(1-a^{2}\right)^{2}(-1+y)(1-\gamma)^{3}(1-\nu)^{3}\Delta^{2}-(1+y)d_{1}^{2}f_{1}^{2}f_{2}^{2}\right\}$
$+4(1+x)(1+x\nu)G(y)\left((1-ab)(\gamma-\nu)(1+\nu)c_{1}-(1-\nu)c_{2}\right){}^{2}g_{5}^{2}$
$+\nu^{-1}(1-\nu)^{3}(\gamma-\nu)\biggr{\{}\left(y^{2}-1\right)G(x)c_{3}^{2}f_{3}^{2}f_{4}^{2}+\left(1-x^{2}\right)G(y)d_{3}^{2}g_{2}^{2}\biggr{\}}$
$+G(x)G(y)\biggr{\{}\frac{x(\gamma-\nu)(c_{1}c_{3}f_{3}f_{6}-bd_{1}f_{1}f_{7})^{2}}{1-\gamma}-(a-b)^{2}y(1-\gamma)(\gamma-\nu)c_{2}^{2}f_{5}^{2}f_{8}^{2}+\frac{(1-a^{2})d_{4}\Delta g_{4}}{\nu}\biggr{\}}\biggr{]},$ (17)
$J(x,y)$ $=\frac{\ell^{2}\tilde{v}_{0}^{2}(1-\gamma)(1+x)(1+y)}{(\gamma^{2}-\nu^{2})(x-y)}\biggr{[}4d_{1}f_{1}f_{2}g_{5}\biggr{\{}(1-ab)(\gamma-\nu)(1+\nu)c_{1}-(1-\nu)c_{2}\biggr{\}}(1+x\nu)(1+y\nu)$
$-(1-x)(1-y)(\gamma-\nu)\biggr{\{}(a-b)(1+x\nu)(1+y\nu)c_{2}f_{5}f_{8}\left(c_{1}c_{3}f_{3}f_{6}-bd_{1}f_{1}f_{7}\right)+(1-\nu)^{3}c_{3}d_{3}f_{3}f_{4}g_{2}\biggr{\}}\biggr{]},$ (18)
$\Omega_{\psi}(x,y)$ $=\frac{\ell\tilde{v}_{0}(1+y)}{H(y,x)}\biggr{[}\nu^{-1}c_{2}f_{5}\left(c_{1}c_{3}f_{4}f_{16}-bd_{1}f_{2}f_{13}\right)(1-x)(1-\nu)(1+x\nu)(1+y\nu)$
$+(1+y\nu)\left\{(1+x)(1+x\nu)c_{1}c_{3}f_{2}f_{9}f_{10}-\frac{1}{2}b\left(1-x^{2}\right)(\gamma-\nu)(1-\nu)c_{3}f_{4}f_{10}f_{12}\right\}$
$-2(1+x)(1-\nu)(1+x\nu)d_{1}f_{2}f_{11}f_{15}\left(c_{3}-(1-ab)(1-\gamma)(1+\nu)\right)$
$+\nu^{-1}c_{3}\left(d_{1}-d_{2}\right)f_{4}g_{3}(1-x)(1-\nu)^{2}(1+x\nu)+c_{3}d_{2}f_{4}g_{10}(1-\nu)^{2}\left(1-x^{2}\right)\biggr{]},$ (19)
$\Omega_{\phi}(x,y)$ $=\frac{\ell\tilde{v}_{0}(1+x)}{H(y,x)}\biggr{[}b(1+x)(1+y)(1+y\nu)d_{1}d_{2}f_{1}f_{13}f_{14}$
$+\frac{(1+x)\left(1-y^{2}\right)(1-\nu)\nu c_{3}\left(b(1+\nu)d_{1}-2(a-b)(1-\gamma)^{2}(1-\nu)\nu\right)f_{3}g_{11}}{1+\nu}$
$+\frac{(y-1)(1-\gamma)(1-\nu)\nu(1+y\nu+x(1+y(4-3\nu)))c_{3}f_{1}f_{3}f_{10}}{1+\nu}$
$+\frac{2(a-b)(1-\gamma)^{2}\left((y-1)(x+y)(1-\nu)^{3}\nu c_{3}f_{3}g_{9}+2(1+x\nu)(1+y\nu)^{2}d_{1}f_{1}g_{1}\right)}{1+\nu}\biggr{]},$ (20)
where
$\begin{split}&\tilde{v}_{0}:=\frac{v_{0}}{\sqrt{\Delta}}:=\sqrt{\frac{2(\gamma^{2}-\nu^{2})}{(1-a^{2})(1-\gamma)\Delta}},\\ &c_{1}:=(1-\gamma)a+(\gamma-\nu)b,\\ &c_{2}:=2a(1-\gamma)\nu+b(\gamma-\nu)(1+\nu),\\ &c_{3}:=2(1-\gamma)\nu+b^{2}(\gamma-\nu)(1+\nu),\\ &d_{1}:=(\nu+1)c_{1}^{2}-(1-\gamma)(1-\nu)^{2},\\ &d_{2}:=b(\nu+1)c_{1}(\gamma-\nu)+2\nu(1-\gamma)(1-\nu),\\ &d_{3}:=\left(a^{2}-1\right)b(\gamma-1)(\nu+1)-ac_{3},\\ &d_{4}:=b^{2}(\gamma-\nu)\left[(\nu+1)^{2}c_{1}^{2}\left(-3(1-\gamma)\nu-\nu^{2}+1\right)-(1-\gamma)(1-\nu)^{4}(2\nu+1)\right]\\ &\quad+(1-\gamma)\left[\left((1-\nu)c_{2}-2\nu^{2}c_{1}\right)^{2}-4\nu^{2}c_{1}^{2}\left(-\gamma(\nu+2)+3\nu^{2}+1\right)\right],\\ &d_{5}:=(1-\gamma)(1-\nu)^{3}[(\gamma-3\nu)\left(b^{2}(\nu-1)(\gamma-\nu)-c_{1}^{2}\right)-2bc_{1}(3\nu-1)(\gamma-\nu)],\\ &d_{6}:=c_{1}(1-\gamma)(1-\nu)\left(1-\nu^{2}\right)\left(c_{2}-(1-\gamma)(a-b)(\gamma-\nu)\right),\\ &\Delta:=f_{8}^{2}-\frac{d_{1}(f_{1}^{2}-f_{8}^{2})}{(1-a^{2})(1-\gamma)(1-\nu^{2})}+\frac{c_{3}(f_{3}^{2}-f_{8}^{2})}{(1-a^{2})(1-\gamma)(1+\nu)},\end{split}$ (21)
and the metric depends on the parameter $\beta$ induced by the Ehlers transformation through coefficients $\{f_{i}\}_{i=1,\dots 16}$ and $\{g_{i}\}_{i=1,\dots 11}$, which are quadratic and quartic polynomials of $\beta$ presented in Appendix A, respectively.
We assume that the coordinates $(t,\psi,\phi,x,y)$ run in the ranges,
$-\infty<t<\infty,\quad 0\leq\psi\leq 2\pi,\quad 0\leq\phi\leq 2\pi,$ (22)
and
$-1\leq x\leq 1,\quad-1/\nu\leq y\leq-1.$ (23)
The solution is described by seven parameters $(\ell,\nu,\gamma,a,b,\beta,\alpha)$, among which the parameters $\ell,\nu,\gamma$, are assumed to lie within the following range
$\ell>0,\quad 0<\nu<\gamma<1,$ (24)
while the ranges for $a,b,\beta,\alpha$ are not specified explicitly at this point. Moreover $\tilde{v}_{0}^{2}$ must be real so that the metric functions take real values, which leads to
$(1-a^{2})\Delta>0.$ (25)
III Boundary conditions
In order that the solution (10) and (II.2) describes a capped black hole of physical interest, we must impose the appropriate boundary conditions at
  • (i)
    infinity $\partial\Sigma_{\infty}=\{(x,y)|x\to y\to-1\}$,
  • (ii)
    the $\phi$-rotational axis $\partial\Sigma_{\phi}=\{(x,y)|x=-1,-1/\nu<y<-1\}$,
  • (iii)
    the $\psi$-rotational axis $\partial\Sigma_{\psi}=\{(x,y)|-1<x<1,y=-1\}$,
  • (iv)
    the inner axis (bubble) $\partial\Sigma_{\rm in}=\{(x,y)|x=1,-1/\nu<y<-1\}$,
  • (vi)
    the center at $(x,y)=(1,-1)$,
  • (v)
    the horizon $\partial\Sigma_{\cal H}=\{(x,y)|-1<x<1,y=-1/\nu\}$.
The boundary condition at infinity (i) is imposed so that the spacetime is asymptotically flat. We require that on the axes (ii), (iii) and (iv), there appear no Dirac-Misner strings, and orbifold singularities at isolated points must be eliminated. We also demand that the horizon (v) describes a smooth null surface whose spatial cross section has a topology of $S^{3}$. At these boundaries, the spacetime is required to allow neither CTCs nor (conical and curvature) singularities. Since the Kretchman invariant can be written as $R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}\sim H(x,y)^{-6}D^{-6}$, the metric (10) exhibits the curvature singularity if there exist the surface $H(x,y)=0$ or $D=0$ on and outside the horizon. Henceforth, we must choose the parameters suitably such that $H(x,y)\neq 0$ and $D\neq 0$ in the region (23) including the boundary. Since we observe from Eqs. (24) and (25) that at infinity,
$H(x=-1,y=-1)=8(1-\gamma)^{3}(1-\nu)^{4}(1-a^{2})\Delta>0,\quad D(x=-1,y=-1)=1>0,$ (26)
the requirements $H(x,y)\neq 0$ and $D\neq 0$ on and outside the horizon can be replaced with $H(x,y)>0$ and $D>0$,
III.1 Asymptotic infinity ($x\to y\to-1$)
The boundary $\partial\Sigma_{\infty}=\{(x,y)|x\to y\to-1\}$ corresponds to the asymptotic infinity. This can be seen in the $(r,\theta)$ coordinates introduced by
$x=-1+\frac{4\ell^{2}(1-\nu)\cos^{2}\theta}{r^{2}},\quad y=-1-\frac{4\ell^{2}(1-\nu)\sin^{2}\theta}{r^{2}},$ (27)
in which the metric (10) asymptotes to the five-dimensional Minkowski metric at the limit $r\to\infty$
$ds^{2}$ $\simeq-\left(1-\frac{8G_{5}M}{3\pi r^{2}}\right)dt^{2}-\frac{8G_{5}J_{\psi}\sin^{2}\theta}{\pi r^{2}}dtd\psi-\frac{8G_{5}J_{\phi}\cos^{2}\theta}{\pi r^{2}}dtd\phi$
$+dr^{2}+r^{2}\sin^{2}\theta d\psi^{2}+r^{2}\cos^{2}\theta d\phi^{2}+r^{2}d\theta^{2},$ (28)
where $M,J_{\psi},J_{\phi}$ are the ADM mass and angular momenta which are expressed as
$M$ $=(c^{2}+s^{2})\frac{3\pi\ell^{2}\tilde{v}_{0}^{2}\left(d_{1}-(1-\nu)c_{3}\right)}{8G_{5}(1-\gamma)(1+\nu)(\gamma-\nu)}\biggr{[}f_{8}^{2}+\frac{(1+\nu)c_{2}^{2}\left(f_{8}^{2}-f_{5}^{2}\right)}{(1-\nu)(\gamma+\nu)\left(d_{1}-(1-\nu)c_{3}\right)}$
$\quad+\frac{(1+\nu)d_{1}\left((1-\gamma)\left(f_{8}^{2}-f_{1}^{2}\right)-(1+\nu)\left(f_{8}^{2}-f_{2}^{2}\right)\right)}{(1-\nu)(\gamma+\nu)\left(d_{1}-(1-\nu)c_{3}\right)}\biggr{]},$ (29)
$J_{\psi}$ $=c^{3}J^{\rm vac}_{\psi}+s^{3}J^{\rm vac}_{\phi},$ (30)
$J_{\phi}$ $=c^{3}J^{\rm vac}_{\phi}+s^{3}J^{\rm vac}_{\psi},$ (31)
with
$J^{\rm vac}_{\psi}$ $=-\frac{\pi\ell^{3}\tilde{v}^{3}_{0}\left(c_{3}f_{4}g_{3}\left(d_{1}-d_{2}\right)+c_{2}f_{5}\left(c_{1}c_{3}f_{4}f_{16}-bd_{1}f_{2}f_{13}\right)\right)}{8G_{5}\nu(1-\gamma)^{2}(\gamma^{2}-\nu^{2})},$ (32)
$J^{\rm vac}_{\phi}$ $=-\frac{\pi\ell^{3}\tilde{v}^{3}_{0}\left(2\nu c_{3}f_{3}f_{1}f_{10}-(2\nu c_{3}f_{3}g_{9}+d_{1}f_{1}g_{1})(a-b)(1-\gamma)(1-\nu)\right)}{4G_{5}(1-\gamma)(1-\nu^{2})(\gamma^{2}-\nu^{2})}.$ (33)
From the asymptotic behavior of the gauge potential (II.2), one can determine the electric charge
$Q$ $:=\frac{1}{8\pi G_{5}}\int_{S}\left(\star F+\frac{1}{\sqrt{3}}F\wedge A\right)$
$=\frac{1}{8\pi G_{5}}\int_{S_{\infty}}\star F$
$=-\frac{2M\tanh(2\alpha)}{\sqrt{3}},$ (34)
where $S$ is an $S^{3}$-surface that encloses both $\partial\Sigma_{\cal H}$ and $\partial\Sigma_{\rm in}$, and $S_{\infty}$ represents spatial infinity. Note that, in the second line, we dropped the Chern-Simons term as it decays faster than the first term at the infinity.
III.2 Rotational Axes ($y=-1,x=\pm 1$)
The boundaries $\partial\Sigma_{\psi}$ and $\partial\Sigma_{\phi}$ are the $\psi$ and $\phi$ rotational axes where the Killing vectors $\partial_{\psi}$ and $\partial_{\phi}$ vanish, respectively. Moreover, on these axes, there exist no conical singularities, due to the periodicities of $\psi$ and $\phi$ specified in Eq. (22).
The inner axis $\partial\Sigma_{\rm in}$ corresponds to the points where the Killing vector $v_{\rm in}=\left(v_{\rm in}^{t},v_{\rm in}^{\psi},1\right)$ vanishes, where
$v_{\rm in}^{t}$ $=\ell\tilde{v}_{0}\left(\frac{c^{3}\tilde{g}_{1}(a-b)}{f_{1}}-\frac{s^{3}g_{12}(1-ab)}{f_{2}}\right),$ (35)
$v_{\rm in}^{\psi}$ $=\frac{\Delta}{f_{1}f_{2}}\left(\frac{ad_{1}+(1-\gamma)(1+\nu)(1-a^{2})c_{1}}{d_{1}}-1\right)+1,$ (36)
with
$\tilde{g}_{1}$ $:=f_{13}f_{14}+\frac{c_{3}f_{3}f_{10}}{2(a-b)(1-\gamma)d_{1}}-\frac{c_{1}c_{3}\left(f_{13}f_{14}-f_{16}f_{19}\right)}{2(a-b)(1-\gamma)^{2}\nu},$ (37)
$g_{12}$ $:=f_{5}f_{13}+\frac{c_{3}(f_{4}f_{17}-f_{5}f_{13})}{2\nu(1-\gamma)(1-ab)}.$ (38)
It is required that there exist no Dirac-Misner string singularities or conical singularities, and the horizon must have the spherical topology of $S^{3}$, which is explained as follows:
  1. 1.
    The condition for the absence of Dirac-Misner string singularity: The inner axis $\partial\Sigma_{\mathrm{in}}$ is free from Dirac-Misner string singularities if the Killing vector $v_{\mathrm{in}}$ has no time component. From Eq. (35), this condition connects the parameter $\alpha$ with the other parameters:
    $\tanh^{3}\alpha=\frac{(a-b)f_{2}\tilde{g}_{1}}{(1-ab)f_{1}g_{12}},$ (39)
    where the magnitude of the right-hand side must be less than or equal to one.
  2. 2.
    The condition for the horizon topology: The cross-section of the Killing horizon has the topology of $L(n;1)$ ($n$: integer) if
    $v_{\rm in}^{\psi}$ $=\frac{\Delta}{f_{1}f_{2}}\left(\frac{ad_{1}+(1-\gamma)(1+\nu)(1-a^{2})c_{1}}{d_{1}}-1\right)+1=n.$ (40)
    To specifically ensure a topology of $S^{3}$, we must set $n=\pm 1$. This work focuses on $n=1$.
  3. 3.
    The condition for the absence of conical singularities: Since on this axis, the Killing vector $\partial_{\phi_{+}}:=\partial_{\phi}+\partial_{\psi}$ vanishes under the two conditions previously mentioned, the inner axis does not exhibit a conical singularity if
    $\left(\frac{\Delta\phi_{+}}{2\pi}\right)^{2}=\frac{d_{1}^{2}f_{1}^{2}f_{2}^{2}}{(1-\gamma)^{3}(1-\nu)^{2}(1+\nu)(1-a^{2})^{2}\Delta^{2}}=1.$ (41)
III.2.1 Boundary behavior around $\partial\Sigma_{\psi}$
One can show $\partial\Sigma_{\psi}$ represents a regular rotational axis in the radial coordinate $r$ by $y=-1-C_{0}r^{2}$ with a positive constant $C_{0}$. Then, the metric at $r\to 0$ ($y\to-1$) behaves as
$ds^{2}\simeq\gamma^{0}_{tt}(x)dt^{2}+2\gamma^{0}_{t\phi}(x)dtd\phi+\gamma^{0}_{\phi\phi}(x)d\phi^{2}+\alpha_{0}(x)(dr^{2}+r^{2}d\psi^{2}+G^{-1}(x)dx^{2}),$ (42)
where
$\gamma_{tt}^{0}=-\frac{D|_{y=-1}H(-1,x)}{H(x,-1)},\quad\gamma_{t\phi}^{0}=-\frac{H(-1,x)[c^{3}\Omega_{\phi}(x,-1)-s^{3}\Omega_{\psi}(-1,x)]}{D^{2}|_{y=-1}H(x,-1)},$
$\gamma_{\phi\phi}^{0}=-\frac{D|_{y=-1}F(x,-1)}{H(-1,x)}-\frac{H(-1,x)[c^{3}\Omega_{\phi}(x,-1)-s^{3}\Omega_{\psi}(-1,x)]^{2}}{D^{2}|_{y=-1}H(x,-1)},\quad\alpha_{0}=\frac{C_{0}\ell^{2}D|_{y=-1}H(x,-1)}{2(1-\gamma)^{3}(1-\nu)^{3}\Delta(1-a^{2})(1+x)^{2}}.$ (43)
Under the assumptions of Eq. (25), $H(x,y)>0$ and $D>0$, one can show that $\alpha_{0}>0$ and
${\rm det}\,(\gamma^{0})=-\frac{16\ell^{2}(1-\gamma)^{3}(1-\nu)^{4}(1-a^{2})\Delta(1-x)(1+\nu x)}{(1+x)D|_{y=-1}H(x,-1)}<0,$ (44)
meaning that $\gamma^{0}_{AB}(A,B=t,\phi)$ is a nonsingular and non-degenerate matrix for $-1<x<1$. Therefore, the metric is regular at $y=-1$.
III.2.2 Boundary behavior around $\partial\Sigma_{\phi}$ and $\partial\Sigma_{\rm in}$
Similarly, $\partial\Sigma_{\phi}$ and $\partial\Sigma_{\rm in}$ can be shown to be the $\phi$-rotational axis and inner rotational axes, respectively. Using the radial coordinate $r$ by $x=\pm 1\mp C_{\pm}r^{2}$ with positive constants $C_{\pm}$ for $x=\pm 1$, we can see that, with the use of Eqs. (39), (40) and (41), the metric at $r\to 0$ ($x\to\pm 1$) behaves as
$ds^{2}\simeq\gamma^{\pm}_{tt}(y)dt^{2}+2\gamma^{\pm}_{t\psi}(y)dtd\psi_{\pm}+\gamma^{\pm}_{\psi\psi}(y)d\psi_{\pm}^{2}+\alpha_{\pm}(y)(dr^{2}+r^{2}d\phi_{\pm}^{2}-G^{-1}(y)dy^{2}),$ (45)
where
$\gamma_{tt}^{\pm}=-\frac{D|_{x=\pm 1}H(y,\pm 1)}{H(\pm 1,y)},\quad\gamma_{t\psi}^{\pm}=-\frac{H(y,\pm 1)[c^{3}\Omega_{\phi}(\pm 1,y)-s^{3}\Omega_{\phi}(y,\pm 1)]}{D^{2}|_{x=\pm 1}H(\pm 1,y)},$
$\gamma_{\psi\psi}^{\pm}=\frac{D|_{x=\pm 1}F(y,\pm 1)}{H(y,\pm 1)}-\frac{H(y,\pm 1)[c^{3}\Omega_{\psi}(\pm 1,y)-s^{3}\Omega_{\phi}(y,\pm 1)]^{2}}{D^{2}|_{x=\pm}H(\pm 1,y)},$
$\alpha_{\pm}=\frac{C_{\pm}\ell^{2}D|_{x=\pm 1}H(\pm 1,y)}{2(1-\gamma)^{3}(1\pm\nu)(1-\nu)^{2}\Delta(1-a^{2})(1\mp y)^{2}},$ (46)
and
$(\psi_{-},\phi_{-}):=(\psi,\phi),\ (\psi_{+},\phi_{+}):=(\psi-\phi,\phi).$ (47)
Under the assumptions $H(x,y)>0$ and $D>0$, we can also show that $\alpha_{\pm}>0$, and
${\rm det}\,(\gamma^{+})=\frac{16\ell^{2}d_{1}^{2}(1+\nu)(1+y)(1+\nu y)}{(1-a^{2})(1-y)D|_{x=1}H(1,y)}<0,$ (48)
${\rm det}\,(\gamma^{-})=\frac{16\ell^{2}(1-\gamma)^{3}(1-\nu)^{4}(1-a^{2})(1-y)(1+\nu y)}{(1+y)D|_{x=-1}H(-1,y)}<0,$ (49)
hence $\gamma^{\pm}_{AB}(A,B=t,\psi)$ is a nonsingular and non-degenerate matrix for $-1/\nu<y<-1$. Thus, the metric is regular at $x=\pm 1$.
Each point in $\partial\Sigma_{\rm in}$ has the topology of $S^{1}$ generated by $\psi_{+}$, which shrinks to a point at $y=-1$ but does not at $y=-1/\nu$. Therefore, $\partial\Sigma_{\rm in}$ has the topology of two-dimensional disk, and called disk-shaped bubble. One can define the two-dimensional area of the bubble ${\mathcal{D}}$ by
[…]
III.3 The center $(x,y)=(1,-1)$
At $(x,y)=(1,-1)$ where the $\psi$-rotational axis and inner rotational axis intersect, the conditions $H(x,y)>0$ and $D>0$ lead to further inequalities
$H(x=1,y=-1)=-8(1-\gamma)(1-\nu)^{2}(1+\nu)d_{1}f_{2}^{2}>0\quad\Longleftrightarrow\quad d_{1}=(\nu+1)c_{1}^{2}-(1-\gamma)(1-\nu)^{2}<0,$ (51)
$D(x=1,y=-1)=c^{2}-s^{2}\frac{(1-\gamma)f_{1}^{2}}{(1+\nu)f_{2}^{2}}>0\quad\Longleftrightarrow\quad\tanh^{2}\alpha<\frac{(1+\nu)f_{2}^{2}}{(1-\gamma)f_{1}^{2}}.$ (52)
For later convenience, instead of Eq. (52), we consider the stronger condition independent of $\alpha$
$1<\frac{(1+\nu)f_{2}^{2}}{(1-\gamma)f_{1}^{2}}.$ (53)
Assuming these inequalities and introducing the coordinates $(r,\theta)$ by
$x=1-{\cal C}_{1}(1+\nu)r^{2}\cos^{2}\theta,\quad y=-1-{\cal C}_{1}(1-\nu)r^{2}\sin^{2}\theta,$ (54)
${\cal C}_{1}:=-\frac{(1-\gamma)^{2}(1-a^{2})\Delta}{\ell^{2}d_{1}(c^{2}(1+\nu)f_{2}^{2}-s^{2}(1-\gamma)f_{1}^{2})},$ (55)
one can show that, with the use of Eqs. (39), (40) and (41), the metric approaches at $r\to 0$
$ds^{2}\simeq-dt^{\prime 2}+({\mathcal{C}}_{1}/|{\mathcal{C}}_{1}|))\left[dr^{2}+r^{2}(d\theta^{2}+\sin^{2}\theta d\psi_{+}^{2}+\cos^{2}\theta d\phi_{+}^{2})\right],$ (56)
where $t^{\prime}:=\sqrt{(1-\gamma)(1+\nu)}f_{1}f_{2}t/(c^{2}(1+\nu)f_{2}^{2}-s^{2}(1-\gamma)f_{1}^{2})$ and $(\psi_{+},\phi_{+})=(\psi-\phi,\phi)$. Since the inequalities (25), (51) and (53) guarantees ${\mathcal{C}}_{1}>0$, the spacetime metric becomes the Minkowski metric. Therefore, the point $(x,y)=(1,-1)$ turns out to be a regular center of the Minkowski spacetime.
III.4 Horizon
The boundary $\partial\Sigma_{\cal H}=\{(x,y)|-1<x<1,y=-1/\nu\}$ corresponds to a Killing horizon with the surface gravity
$\kappa=\frac{2\sqrt{2}\sqrt{1-\gamma}(\gamma-\nu)^{3/2}\sqrt{\nu(1+\nu)}(\gamma+\nu)}{\ell c_{3}\tilde{v}_{0}^{3}\left(c^{3}\left(ac_{1}-1+\nu\right)f_{3}g_{3}+(a-b)s^{3}(\gamma-\nu)f_{4}g_{13}\right)},$ (57)
and the horizon angular velocities $v_{\cal H}=(1,\omega_{\psi},\omega_{\phi})$:
$\omega_{\psi}:=-\frac{(\gamma-\nu)f_{3}f_{4}}{\ell\tilde{v}_{0}\left(c^{3}\left(ac_{1}-1+\nu\right)f_{3}g_{3}+s^{3}(a-b)(\gamma-\nu)f_{4}g_{13}\right)},$ (58)
$\omega_{\phi}:=\frac{(\gamma-\nu)d_{3}g_{2}}{\ell\tilde{v}_{0}c_{3}\left(c^{3}\left(ac_{1}-1+\nu\right)f_{3}g_{3}+s^{3}(a-b)(\gamma-\nu)f_{4}g_{13}\right)},$ (59)
where
$g_{13}:=f_{8}f_{18}-\frac{f_{1}f_{10}}{2(1-\gamma)(1-\nu)(b-a)}.$ (60)
The spatial cross section of the horizon has topology $S^{3}$ under the boundary condition (40) with $n=1$. The area of the horizon cross section at $y=-1/\nu$ is given in terms of the surface gravity
$A_{\cal H}=8\pi^{2}\ell^{2}\nu\kappa^{-1}.$ (61)
Let us show that both the metric and gauge potential are regular at $y=-1/\nu$, by introducing the ingoing/outgoing Eddington-Finkelstein coordinates by
$dx^{i}=dx^{\prime i}\pm v_{\cal H}^{i}\frac{(1-\nu^{2})}{2\nu\kappa G(y)}dy,$ (62)
where $(x^{i})=(t,\psi,\phi)$ $(i=0,1,2)$ and the metric near $y=-1/\nu$ behaves as
$ds^{2}$ $\simeq\alpha_{H}(x)\left(\frac{4\nu^{2}\kappa^{2}G(y)}{(1-\nu^{2})^{2}}dt^{\prime 2}\pm\frac{4\nu\kappa}{1-\nu^{2}}dt^{\prime}dy+\frac{dx^{2}}{G(x)}\right)$
$+\gamma^{H}_{\psi\psi}(x)(d\psi^{\prime}-\omega_{\psi}dt^{\prime})^{2}+2\gamma^{H}_{\psi\phi}(x)(d\psi^{\prime}-\omega_{\psi}dt^{\prime})(d\phi^{\prime}-\omega_{\phi}dt^{\prime})+\gamma^{H}_{\phi\phi}(x)(d\phi^{\prime}-\omega_{\phi}dt^{\prime})^{2},$ (63)
with
$\begin{split}&\gamma_{\psi\psi}^{H}=\frac{D|_{y=-1/\nu}F(-1/\nu,x)}{H(-1/\nu,x)}-\frac{H(-1/\nu,x)[c^{3}\Omega_{\psi}(x,-1/\nu)-s^{3}\Omega_{\phi}(-1/\nu,x)]^{2}}{D^{2}|_{y=-1/\nu}H(x,-1/\nu)},\\ &\gamma_{\psi\phi}^{H}=-\frac{D|_{y=-1/\nu}J(x,-1/\nu)}{H(-1/\nu,x)}-\frac{H(-1/\nu,x)[c^{3}\Omega_{\psi}(x,-1/\nu)-s^{3}\Omega_{\phi}(-1/\nu,x)][c^{3}\Omega_{\phi}(x,-1/\nu)-s^{3}\Omega_{\psi}(-1/\nu,x)]}{D^{2}|_{y=-1/\nu}H(x,-1/\nu)},\\ &\gamma_{\phi\phi}^{H}=-\frac{D|_{y=-1/\nu}F(x,-1/\nu)}{H(-1/\nu,x)}-\frac{H(-1/\nu,x)[c^{3}\Omega_{\phi}(x,-1/\nu)-s^{3}\Omega_{\psi}(-1/\nu,x)]^{2}}{D^{2}|_{y=-1/\nu}H(x,-1/\nu)},\\ &\alpha_{H}=\frac{\ell^{2}\nu^{2}D|_{y=-1/\nu}H\left(x,-1/\nu\right)}{4(1-\gamma)^{3}(1-\nu)^{2}\left(1-a^{2}\right)\Delta(1+\nu x)^{2}}.\end{split}$ (64)
Hence, under the assumptions $H(x,y)>0$ and $D>0$, we can show $\alpha_{H}>0$ and
${\rm det}\,(\gamma^{H})=\frac{4\ell^{2}(1-\gamma)^{3}(1-\nu)^{4}(1+\nu)^{2}(1-a^{2})\Delta(1-x^{2})}{\kappa^{2}D(x,y=-1/\nu)H(x,-1/\nu)(1+\nu x)}$ (65)
and thus $\gamma^{H}_{AB}(A,B=\psi,\phi)$ is a nonsingular and non-degenerate matrix for $-1<x<1$, meaning that the metric is smoothly continued to $-\infty<y<-1/\nu$ across the horizon $y=-1/\nu$. Moreover, in the Eddington-Finkelstein coordinate, the gauge potential also remains regular at the horizon $y=-1/\nu$ under the gauge transformation
$A^{\prime}=A\pm d\left(\frac{(1-\nu^{2})\Phi_{e}}{2\nu\kappa}\int\frac{dy}{G(y)}\right),$ (66)
where $\Phi_{e}$ is the electric potential defined by
$\Phi_{e}$ $:=-i_{v_{\cal H}}A\bigr{|}_{y=-1/\nu}=-(A_{t}+\omega_{\psi}A_{\psi}+\omega_{\phi}A_{\phi})\bigr{|}_{y=-1/\nu}$
$=-\frac{\sqrt{3}cs\left(f_{4}g_{13}s(b-a)(\gamma-\nu)-cf_{3}g_{3}\left(ac_{1}+\nu-1\right)\right)}{f_{4}g_{13}s^{3}(b-a)(\gamma-\nu)-c^{3}f_{3}g_{3}\left(ac_{1}+\nu-1\right)}.$ (67)
The points $(x,y)=(\pm 1,-1/\nu)$, at which the rotational $\phi$ or $\phi_{+}(=\phi)$ axes and the horizon intersect, represent merely regular points. By introducing the coordinates $(r,\theta)$ for $(x,y)=(\pm 1,-1/\nu)$ as
$x=\pm 1\mp c_{\pm}r^{2}\sin^{2}\theta,\quad y=-\frac{1}{\nu}\left(1-\frac{(1\mp\nu)c_{\pm}}{2}r^{2}\cos^{2}\theta\right),\quad c_{\pm}:{\rm positive\ constants},$ (68)
and with the use of Eqs. (40) and (41), the metric at $r\to 0$ ($(x,y)\to(\pm 1,-1/\nu)$) behaves as
$ds^{2}\simeq dr^{2}+r^{2}d\theta^{2}+r^{2}\sin^{2}\theta(d\phi_{\pm}-\omega_{\phi}dt)^{2}-r^{2}\kappa^{2}\cos^{2}\theta dt^{2}+R_{\pm}^{2}(d\psi_{\pm}-\omega_{\psi}^{\pm}dt)^{2},$ (69)
where $\omega_{\psi}^{+}=\omega_{\psi}-\omega_{\phi}$ , $\omega_{\psi}^{-}=\omega_{\psi}$ , and $(\psi_{\pm},\phi_{\pm})$ are defined in Eq. (47). $R_{\pm}:=\sqrt{g_{\psi\psi}}|_{(x,y)=(\pm 1,-1/\nu)}$ are given by
$R_{+}=-\frac{2d_{1}\ell(\gamma-\nu)\sqrt{\nu(\gamma+\nu)}\left(bf_{1}f_{13}f_{15}s^{3}\left(d_{2}-(1-\gamma)(1-\nu)(\gamma+\nu)\right)+c^{3}d_{2}f_{2}f_{11}f_{14}\right)}{\sqrt{\left(1-a^{2}\right)(1-\gamma)\Delta}\left(-c^{2}d_{2}^{2}f_{14}^{2}(\gamma-\nu)-2f_{15}^{2}\nu s^{2}\left(d_{2}-(1-\gamma)(1-\nu)(\gamma+\nu)\right){}^{2}\right)},$ (70)
$R_{-}=\frac{2\ell\sqrt{\nu(\nu+1)(\gamma+\nu)}\left(f_{4}g_{13}s^{3}(b-a)(\gamma-\nu)+c^{3}f_{3}g_{3}\left(-ac_{1}-\nu+1\right)\right)}{\sqrt{\left(1-a^{2}\right)\Delta}\left(2c^{2}(1-\gamma)f_{3}^{2}\nu+f_{4}^{2}(\nu+1)s^{2}(\gamma-\nu)\right)},$ (71)
and we also set $c_{\pm}$ as
$c_{\pm}=\frac{(1\pm\nu)R_{\pm}\kappa}{2\ell^{2}\nu(1\mp\nu)}.$ (72)
In the Cartesian coordinates $(T,X,Y,Z,W)=(\kappa t,r\cos\theta,r\sin\theta\cos(\phi_{\pm}-\omega_{\phi}t),r\sin\theta\sin(\phi_{\pm}-\omega_{\phi}t),R_{\pm}(\psi_{\pm}-\omega^{\pm}_{\psi}t))$, the above asymptotic metric becomes
$ds^{2}\simeq-X^{2}dT^{2}+dX^{2}+dY^{2}+dZ^{2}+dW^{2},$ (73)
where the Rindler horizon lies at $X=0$. Therefore, the metric is regular at $(x,y)=(\pm 1,-1/\nu)$.
III.5 Absence of curvature singularities and CTCs
In the previous subsections, we have derived the conditions ensuring the regularity of the solution at the boundaries of the C-metric coordinates $(x,y)$. Here, we demonstrate that these conditions are also sufficient to guarantee the solution’s regularity throughout the entire region spanned by the C-metric coordinates, as specified in Eq. (23). We show that under the inequalities (25), (51) and (53), the conditions (39), (40) and (41) ensures $H(x,y)>0$ and $D>0$ within the region specified in (23).
Figure 1: Profile of $H(x,y)$ for $(x,y)\in[-1,1]\times[-1/\nu,-1]$ and $(\sigma,\tilde{\nu},\tilde{\gamma})=(0.8,1.2,0.239)$.
It is challenging to demonstrate analytically $H(x,y)>0$, however, as depicted in Fig. 1, we can numerically confirm $H(x,y)>0$ for the parameters satisfying the conditions (39), (40) and (41), as well as the inequalities (25), (51) and (53). Once $H(x,y)>0$ is established, it follows that $D>0$ within the ranges (23) because of the relation $D=1+s^{2}(H(x,y)-H(y,x))/H(x,y)\geq 1$ and the inequality
$H(x,y)-H(y,x)$ $=(x-y)\biggr{[}\frac{(1-\nu)\left((\gamma-\nu)d_{2}^{2}f_{14}^{2}+2\nu\left(d_{2}-(1-\gamma)(1-\nu)(\gamma+\nu)\right){}^{2}f_{15}^{2}\right)}{\gamma-\nu}(1+x)(-1-y)$
$+\frac{4(1-\gamma)(1-\nu)(-d_{1})(1+\nu)f_{2}^{2}}{1+\nu}\left(1-\frac{(1-\gamma)f_{1}^{2}}{(1+\nu)f_{2}^{2}}\right)(1+\nu x)(1+\nu y)$
$+\frac{(1-\gamma)(1-\nu)^{3}c_{3}\left(2(1-\gamma)\nu f_{3}^{2}+(\gamma-\nu)(1+\nu)f_{4}^{2}\right)}{1+\nu}(1-x)(1-y)\biggr{]}\geq 0,$ (74)
where nonnegativity of the second term follows from Eq. (53).
The absence of closed timelike curves (CTCs) is demonstrated if the two-dimensional metric $g_{IJ}(I,J=\psi,\phi)$ is positive definite, except at the axes where $x=\pm 1$ and $y=-1$. This corresponds to the conditions $\mathrm{det}(g_{IJ})>0$ and $\mathrm{tr}(g_{IJ})>0$ for $-1<x<1$ and $-1/\nu\leq y<-1$. Following the reasoning discussed in Ref. Suzuki:2024coe , demonstrating that $\mathrm{det}(g_{IJ})>0$ is sufficient.
$\mathrm{det}(g_{IJ})$ vanishes on the $\phi$ and $\psi$-rotational axes at $x=-1$ and $y=-1$, as well as on the inner axis at $x=1$ after imposing Eq. (39). To demonstrate the positive definiteness of $\mathrm{det}(g_{IJ})$ away from these, we consider the positive definiteness of the alternative function ${\cal D}(x,y)$ for $-1\leq x\leq 1$ and $-1/\nu\leq y\leq-1$, defined by
${\cal D}(x,y):=\frac{(x-y)^{4}DH(x,y)}{\ell^{4}\tilde{v}_{0}^{4}(1-x^{2})(-1-y)}{\rm det}(g_{IJ}).$ (75)
Although giving the analytic proof for the positivity of ${\cal D}(x,y)$ in the entire region is challenging, one can still possible to demonstrate positivity on the horizon:
${\cal D}(x,-1/\nu)=\frac{(1-\gamma)c_{3}^{2}(1-\nu)^{3}(\nu+1)(1+\nu x)^{3}\left[c^{3}f_{3}g_{3}\left(-ac_{1}-\nu+1\right)-f_{4}g_{13}s^{3}(a-b)(\gamma-\nu)\right]{}^{2}}{\nu^{4}(\gamma-\nu)^{2}(\gamma+\nu)}>0.$ (76)
Figure 2: Profile of ${\cal D}(x,y)$ for $(x,y)\in[-1,1]\times[-1/\nu,-1]$ and $(\sigma,\tilde{\nu},\tilde{\gamma})=(0.8,1.2,0.239)$.
For the other region, as shown in Fig. 2, we numerically confirmed the positivity both on and outside the horizon for several choices of the parameters under the constraints (39), (40) and (41) followed by the inequalities (25), (51) and (52).
III.6 Finding parameters for capped black holes
We are now aiming to find a capped black hole solution that does not have any curvature singularities, orbifold singularities, conical singularities, or Dirac-Misner string singularities both on and outside the horizon, additionally, they are free from closed timelike curves. This can be achieved if we can find a set of parameters $(\ell,\nu,\gamma,a,b,\beta,\alpha)$ that meets three constraints (39), (40) and (41). As outlined in our previous works Suzuki:2023nqf ; Suzuki:2024phv , such black holes are distinguished by having a compact inner rotational axis with a two-dimensional disk topology, referred to as a “bubble,” which lies adjacent to the horizon. The existence of this extra bubble implies that the intersection of the DOC and timeslice $\Sigma$ has the nontrivial topology of ${\rm DOC}\cap\Sigma=[{\mathbb{R}}^{4}\#{\mathbb{C}}{\mathbb{P}}^{2}]\setminus{\mathbb{B}}^{4}$, differing it from the Cvetič-Youm black hole with the trivial topology of ${\rm DOC}\cap\Sigma={\mathbb{R}}^{4}\setminus{\mathbb{B}}^{4}$ Cvetic:1996xz .
To solve the constraints, we introduce the new parameters as
$\tilde{\nu}:=\sqrt{1+\nu},\quad\tilde{\gamma}:=\sqrt{1-\gamma},$ (77)
which is helpful in avoiding the square root terms in the later expressions. From Eq. (24), they are restricted to the range
$1<\tilde{\nu}<\sqrt{2},\quad 0<\tilde{\gamma}<\sqrt{2-\tilde{\nu}^{2}}.$ (78)
Then, Eq. (40) can be solved with respect to $a$ as
$a=\frac{\tilde{\gamma}^{2}\left(\tilde{\nu}^{2}-2\right)\left(\left(2b^{2}+b-1\right)\tilde{\nu}^{2}+2\right)+b^{2}\tilde{\nu}^{2}\left(\tilde{\nu}^{2}-2\right)^{2}+b(b+1)\tilde{\gamma}^{4}\tilde{\nu}^{2}}{\tilde{\gamma}^{2}\tilde{\nu}^{2}\left((b+1)\tilde{\gamma}^{2}+b\left(\tilde{\nu}^{2}-2\right)\right)}.$ (79)
Using Eqs. (77) and (79), one can rewrite Eq. (41) as
$\left[\Delta\left(\left(b^{2}-1\right)\tilde{\gamma}^{4}\tilde{\nu}^{2}+\tilde{\gamma}^{2}\left(\tilde{\nu}^{2}-2\right)\left(\left(2b^{2}-1\right)\tilde{\nu}^{2}+2\right)+b^{2}\tilde{\nu}^{2}\left(\tilde{\nu}^{2}-2\right)^{2}\right)+f_{1}f_{2}\tilde{\gamma}^{3}\tilde{\nu}\left(\tilde{\nu}^{2}-2\right)\right]$
×$\times\left[\Delta\left(\left(b^{2}-1\right)\tilde{\gamma}^{4}\tilde{\nu}^{2}+\tilde{\gamma}^{2}\left(\tilde{\nu}^{2}-2\right)\left(\left(2b^{2}-1\right)\tilde{\nu}^{2}+2\right)+b^{2}\tilde{\nu}^{2}\left(\tilde{\nu}^{2}-2\right)^{2}\right)-f_{1}f_{2}\tilde{\gamma}^{3}\tilde{\nu}\left(\tilde{\nu}^{2}-2\right)\right]=0.$ (80)
Since, with the use of Eq. (79), $f_{1},f_{2}$ and $\Delta$ is invariant under the transformation $\tilde{\gamma}\to-\tilde{\gamma}$, Eq. (80) reduces to a single condition by letting $\tilde{\gamma}$ to take negative values
$\Delta\left(\left(b^{2}-1\right)\tilde{\gamma}^{4}\tilde{\nu}^{2}+\tilde{\gamma}^{2}\left(\tilde{\nu}^{2}-2\right)\left(\left(2b^{2}-1\right)\tilde{\nu}^{2}+2\right)+b^{2}\tilde{\nu}^{2}\left(\tilde{\nu}^{2}-2\right)^{2}\right)+f_{1}f_{2}\tilde{\gamma}^{3}\tilde{\nu}\left(\tilde{\nu}^{2}-2\right)=0.$ (81)
Moreover, with Eq. (79), the condition (51) is simplified as
$(d_{1}=)-\tilde{\gamma}^{2}(2-\tilde{\nu}^{2})^{2}(1-\tilde{b}^{2})<0\quad\Longleftrightarrow\quad-1<\tilde{b}<1,$ (82)
where another new parameter $\tilde{b}$ is introduced by
$b=-\frac{\tilde{\gamma}\left(\tilde{b}\tilde{\gamma}\tilde{\nu}-\tilde{\nu}^{2}+2\right)}{\tilde{b}\tilde{\nu}\left(\tilde{\gamma}^{2}+\tilde{\nu}^{2}-2\right)}.$ (83)
By substituting Eq. (83), Eqs. (79) and (81) can be written in terms of the new parameters $(\tilde{\nu},\tilde{\gamma},\tilde{b})$ as
$\Delta\left((\tilde{b}^{2}-1)\left(2-\tilde{\nu}^{2}\right)-2\tilde{b}\tilde{\gamma}\tilde{\nu}\right)+f_{1}f_{2}\tilde{b}^{2}\tilde{\gamma}\tilde{\nu}=0,$ (84)
$a=\frac{(1-\tilde{b}^{2})\left(\tilde{\nu}^{2}-2\right)-\tilde{b}\tilde{\gamma}\tilde{\nu}}{\tilde{b}\tilde{\gamma}\tilde{\nu}}.$ (85)
Additionally, to simplify the expression, we replace the parameter $\alpha$ with
$\sigma:=\tanh\alpha,\quad-1<\sigma<1,$ (86)
where Eq. (39) is written as
$\sigma^{3}=\frac{\tilde{b}\tilde{\nu}\left(\tilde{b}\tilde{\gamma}\tilde{\nu}+\tilde{b}^{2}\left(\tilde{\gamma}^{2}-2\right)+\left(\tilde{b}^{2}-1\right)\tilde{\nu}^{2}+2\right)}{\tilde{\gamma}\left(\tilde{b}\left(\tilde{b}^{2}-2\right)\tilde{\gamma}\tilde{\nu}+\left(1-2\tilde{b}^{2}\right)\tilde{\nu}^{2}+2\left(\tilde{b}^{2}-1\right)\right)}\frac{f_{2}\tilde{g}_{1}}{f_{1}g_{12}}.$ (87)
Note that, from Eq. (III.1), $\sigma$ is directly related to the charge-to-mass ratio as
$\frac{Q}{M}=-\frac{4\sigma}{\sqrt{3}(1+\sigma^{2})},$ (88)
where, in particular, the limit $\sigma\to\pm 1(\alpha\to\pm\infty)$ corresponds to the BPS limit $|Q|/M\to 2/\sqrt{3}$.
Therefore, to obtain a physical solution, one must solve Eqs. (84) and (87) under the conditions (25) and (52) for the parameters $(\tilde{\gamma},\tilde{\nu},\tilde{b},\sigma,\beta)$ within the specified range
$1\displaystyle 1<\tilde{\nu}<\sqrt{2},\quad 0<|\tilde{\gamma}|<\sqrt{2-\tilde{\nu}^{2}},\quad-1<\tilde{b}<1,\quad-1<\sigma<$ 1 . (89)
Since Eqs. (84) and (87) yield quartic and sixth-order equations of $\beta$, respectively, finding the analytic solution with respect to $\beta$ is not easy. Therefore, we resort to exploring the solutions numerically. Since $\beta$ and $\tilde{b}$ are represented as the functions of $\sigma,\tilde{\nu}$ and $\tilde{\gamma}$ through Eqs. (84) and (87), the physical solution is uniquely determined by a given $(\sigma,\tilde{\nu},\tilde{\gamma})$ in the range (89). In Fig. 3, we illustrate all possible solution for $\sigma=0.8,\ \tilde{\nu}=1.2$ and $0<|\tilde{\gamma}|<\sqrt{2-\tilde{\nu}^{2}}$. Moreover, we find that the metric and gauge potential is invariant under the transformation $\tilde{\gamma}\to-\tilde{\gamma}$, despite $\beta,\tilde{b}$ themselves not being invariant [^1] The metric function $H(x,y),F(x,y),J(x,y)$ are modified by the transformation $\tilde{\gamma}\to-\tilde{\gamma}$ but only upto a common factor $H(x,y)\to CH(x,y),\ F(x,y)\to CF(x,y),\ J(x,y)\to CJ(x,y)$. . As in the black ring case Suzuki:2024vzq , this might imply the possible extension from the $\tilde{\gamma}=0$ phase into the parameter region $\gamma\ (=1-\tilde{\gamma}^{2})>1$ where $\tilde{\gamma}$ becomes pure imaginary and $\beta,\tilde{b}$ become complex. In this paper, however, we restrict ourselves for the phases with all the parameters real.
Figure 3: All solutions for $\sigma=0.8$ and $\tilde{\nu}=1.2$ shown in the $(\tilde{\gamma},\beta)$ and $(\tilde{\gamma},\tilde{b})$ plains. $\beta$ diverges at $\tilde{\gamma}_{\infty}=-0.195...$ and the spacetime seems to be singular there. But the phases are smoothly continued at $\tilde{\gamma}=\tilde{\gamma}_{\infty}$ if we use the other parameter $\beta^{\prime}=\beta^{-1}$ around there. Despite $\beta=\beta(\tilde{\gamma})$ and $\tilde{b}=\tilde{b}(\tilde{\gamma})$ do not exhibit the mirror symmetry with respect to $\tilde{\gamma}$, the metric and gauge functions coincide for $\tilde{\gamma}=C>0$ (blue curve) and $\tilde{\gamma}=-C$ (red curve).
IV Physical properties
Finally, we study the phase space of the new solution. The new capped black hole solution is determined by four independent parameters $(\ell,\sigma,\tilde{\nu},\tilde{\gamma})$. Here, we focus on the phase diagram in the $\sigma=0.8$ case. Other cases are qualitatively the same.
IV.1 Thermodynamics
From Ref. Kunduri:2013vka , the capped black hole follows the first law, which involves local quantities on the bubble ${\mathcal{D}}$
$\delta M=\frac{\kappa}{8\pi}\delta A_{\cal H}+V_{\psi}\delta J_{\psi}+V_{\phi}\delta J_{\phi}+\frac{1}{2}\Phi_{\cal H}\delta Q+{\cal Q}_{\mathcal{D}}\delta\Phi_{\mathcal{D}},$ (90)
and the Smarr formula
$M=\frac{3\kappa A_{\cal H}}{16\pi}+\frac{3}{2}V_{\psi}J_{\psi}+\frac{3}{2}V_{\phi}J_{\phi}+\frac{1}{2}\Phi_{\cal H}Q+\frac{1}{2}{\cal Q}_{\mathcal{D}}\Phi_{\mathcal{D}},$ (91)
where $\Phi_{\cal H}$ is the electric potential (67), $\Phi_{\mathcal{D}}$ and ${\cal Q}_{\mathcal{D}}$ are the magnetic potential and magnetic flux on ${\mathcal{D}}$ defined as
$\Phi_{\mathcal{D}}$ $:=-i_{v_{\rm in}}A\biggr{|}_{x=1}=-(A_{\phi}+A_{\psi})\biggr{|}_{x=1}=-\frac{\sqrt{3}\ell cs\tilde{v}_{0}\left(2c(1-\gamma)f_{2}\nu\tilde{g}_{1}(a-b)-f_{1}s\left(c_{3}f_{4}f_{17}-bc_{2}f_{5}f_{13}\right)\right)}{2\nu(1-\gamma)f_{1}f_{2}},$ (92)
${\cal Q}_{\mathcal{D}}$ $:=\frac{1}{4}\int_{\mathcal{D}}\left[i_{v_{H}}\star F-\frac{1}{\sqrt{3}}(\Phi-\Phi_{\cal H})F\right]=\sqrt{\frac{3}{2}}\frac{\pi\ell scbd_{1}f_{11}f_{13}\sqrt{\left(1-a^{2}\right)\Delta}\sqrt{\gamma^{2}-\nu^{2}}}{2\sqrt{1-\gamma}c_{3}\left(c^{3}f_{3}g_{3}\left(ac_{1}+\nu-1\right)-f_{4}g_{13}s^{3}(b-a)(\gamma-\nu)\right)},$ (93)
with $\Phi(y):=-i_{v_{\cal H}}A|_{x=1}=-(A_{t}+\omega_{\psi}A_{\psi}+\omega_{\phi}A_{\phi})|_{x=1}$. Henceforth, the thermodynamics implies that the capped black hole is parametrized, at most, by the five independent parameters $(M,J_{\psi},J_{\phi},Q,\Phi_{\mathcal{D}})$. However, it is not clear whether these five parameters uniquely determine the solution since the known uniqueness theorem assumes more parameters to specify the solution Armas:2014gga .
Figure 4: The $\sigma=0.8$ phase shown by the colored regions in the $(j_{\psi},j_{\phi})$, $(\phi_{\mathcal{D}},j_{\phi})$ and $(q_{\mathcal{D}},j_{\phi})$ planes. The $\tilde{\nu}={\rm constant}$ phases are shown by thin blue curves with values of $\tilde{\nu}$. To keep the visibility in the panel (a), the $\tilde{\nu}={\rm constant}$ curves are displayed only in the close-up. In each panel, the extremal capped black hole limit, extremal black hole limit and $\tilde{\gamma}=0$ limit are shown by black, orange and black dashed curves, respectively. The red dot-dashed curves denote critical curves for large and small bubble branches, which also become a part of the boundary in the panel (a). In the panel (a), two branches exists for given $(j_{\psi},j_{\phi})$ in the purple region, while only the large bubble branch in the gray region. In the panel (b) and (c), the critical curves separate two branches and the orange dashed curves correspond to the borders between the single and double branch regions. In the panel (b) and (c), the red and blue regions represent to the small bubble branch and corresponding large bubble branch that overlaps in the purple region of the panel (a), respectively, and only the large bubble branch exists in the gray region. In each panel, the extremal capped black hole limit and extremal black hole limit intersect at $A$ and the $\tilde{\gamma}=0$ limit intersects with the former two at $B$ and $C^{\prime}$, respectively. The critical curve has an end on the extremal black hole limit at $E$ and another on the $\tilde{\gamma}=0$ limit at $D$. In the panel (c), the extremal capped black hole limit and $\tilde{\gamma}=0$ limit do not intersect, but instead, are connected by the $\partial_{\tilde{\nu}}j_{\phi}\partial_{\tilde{\gamma}}q_{\mathcal{D}}-\partial_{\tilde{\nu}}q_{\mathcal{D}}\partial_{\tilde{\gamma}}j_{\phi}=0$ curve in the $(q_{\mathcal{D}},j_{\phi})$ plane (blue dot-dashed curve), which intersect with the extremal capped black hole limit at $F$ and meet with the $\tilde{\gamma}=0$ at $C^{\prime}$.
IV.2 Phase diagram for $\sigma=0.8$
To study the phase diagram of the capped black holes, we introduce the dimensionless variables as follows:
$j_{\psi}=\frac{4G_{5}J_{\psi}}{\pi r_{M}^{3}},\quad j_{\phi}=\frac{4G_{5}J_{\phi}}{\pi r_{M}^{3}},\quad a_{\cal H}=\frac{\sqrt{2}A_{\cal H}}{\pi^{2}r_{M}^{3}},\quad\phi_{\mathcal{D}}:=\frac{\Phi_{\mathcal{D}}}{r_{M}},\quad q_{\mathcal{D}}:=\frac{{\cal Q}_{\mathcal{D}}}{r_{M}}.$ (94)
where $r_{M}:=8G_{5}M/(3\pi)$.
In Fig. 4, the phase for $\sigma=0.8$ is depicted in the $(j_{\psi},j_{\phi})$, $(\phi_{\mathcal{D}},j_{\phi})$, and $(q_{\mathcal{D}},j_{\phi})$ planes, where the solution exists within the colored regions. Since the BPS bound is saturated only at the limit $\sigma\to\pm 1$, all the phases for $\sigma=0.8$ are non-supersymmetric. In each panel, curves of constant $\tilde{\nu}$ are shown as thin blue lines. The capped black hole solutions are not uniquely determined by their conserved charges $(M,Q,J_{\psi},J_{\phi})$ and feature two distinct branches, which we call large bubble branch and small bubble branch. As explained in detail below, the former always has the larger bubble than the latter. In Fig. 4(a), both branches coexist within the purple region. These branches are distinguishable by their magnetic flux or potential values. In Figs. 4(b) and (c), the small bubble branch is represented by the red region, while the large bubble branch occupies both blue and gray regions. The portion of the large bubble branch that shares the same $(M,Q,J_{\psi},J_{\phi})$ as the small bubble branch is highlighted in blue. The segment that does not overlap with the small bubble branch in terms of $(M,Q,J_{\psi},J_{\phi})$ is shown in gray.
The phase diagram for $\sigma={\rm constant}$ is bounded by the following phases:
  • Nonsupersymetric extremal black hole limit (orange solid curves in Fig. 4), as $\tilde{\gamma}\to\sqrt{2-\tilde{\nu}^{2}}$ with $\tilde{\nu}$ held constant, where the surface gravity $\kappa$ approaches zero. In this limit, the angular momenta correspond to those of the extremal Cvetič-Youm black holes, maintaining the same charge-to-mass ratio as specified in Eq. (88)
    $j_{\psi}^{\rm BH}=\frac{j_{1}+\sigma^{3}j_{2}}{(1+\sigma^{2})^{3/2}},\quad j_{\phi}^{\rm BH}=\frac{j_{2}+\sigma^{3}j_{1}}{(1+\sigma^{2})^{3/2}},$ (95)
    where
    $j_{1}$ $=1+j_{2}$
    $=\frac{\left(\sigma^{3}+1\right)\left(3\left(\sigma^{3}-3\right)\tilde{\nu}^{2}(\tilde{\nu}^{2}-2)+2\left(\sqrt{\sigma^{6}\left(-3\tilde{\nu}^{4}+6\tilde{\nu}^{2}+1\right)+2\sigma^{3}\left(3\tilde{\nu}^{4}-6\tilde{\nu}^{2}+1\right)+1}-\sigma^{3}-1\right)\right)}{9\tilde{\nu}^{2}\left(2-\tilde{\nu}^{2}\right)}.$ (96)
    The entropy, magnetic flux and potential have the limit
    $a_{\cal H}^{\rm BH}=\frac{2\sqrt{2}\left(\sigma^{3}+1\right)^{3/2}}{9\left(\sigma^{2}+1\right)^{3/2}\tilde{\nu}^{2}\left(2-\tilde{\nu}^{2}\right)}\biggr{[}-9\sigma^{3}\left(\sigma^{6}-5\sigma^{3}+6\right)\tilde{\nu}^{8}+9\sigma^{3}\left(\sigma^{6}-17\sigma^{3}+30\right)\tilde{\nu}^{6}$
    $\quad+6\left(16\sigma^{9}+8\sigma^{6}-83\sigma^{3}-3\right)\tilde{\nu}^{4}-48\left(3\sigma^{9}-4\sigma^{6}-8\sigma^{3}-1\right)\tilde{\nu}^{2}-32\left(\sigma^{3}+1\right)^{3}$
    $\quad+\left(9\sigma^{3}\left(\sigma^{3}-2\right)\tilde{\nu}^{6}+\left(-48\sigma^{6}+96\sigma^{3}+18\right)\tilde{\nu}^{4}+48\left(\sigma^{6}-3\sigma^{3}-1\right)\tilde{\nu}^{2}+32\left(\sigma^{3}+1\right)^{2}\right)$
    ×$\quad\times\sqrt{\sigma^{6}\left(-3\tilde{\nu}^{4}+6\tilde{\nu}^{2}+1\right)+2\sigma^{3}\left(3\tilde{\nu}^{4}-6\tilde{\nu}^{2}+1\right)+1}\biggr{]}^{1/2},$ (97)
    $\phi_{\mathcal{D}}^{\rm BH}=\frac{\sqrt{3}\sigma^{2}}{\sqrt{1+\sigma^{2}}},\quad q_{\mathcal{D}}^{\rm BH}=0.$ (98)
    While, the entropy of the Cvetič-Youm black hole, in general, is given by Cvetic:1996kv
    $a_{\cal H}^{\rm CY}=\frac{\sqrt{2}\left(\sqrt{1-\left(j_{1}-j_{2}\right){}^{2}}\left(1-\sigma^{3}\right)+\sqrt{1-\left(j_{1}+j_{2}\right){}^{2}}\left(1+\sigma^{3}\right)\right)}{\left(1+\sigma^{2}\right)^{3/2}},$ (99)
    with $(j_{1},j_{2})$ in Eq. (96). Since the Cvetič-Youm black hole does not possess the bubble of $S^{2}$ outside the horizon, both the magnetic flux and magnetic potential are trivially zero. We note that the limit of the entropy (97) does not coincide with that of of the corresponding Cvetič-Youm black hole in Eq. (99). The similar discontinuity is reported for the black hole limit in the black ring phase Elvang:2007hs ; Elvang:2004rt . As speculated in Ref. Elvang:2004rt , this may be because the spacetime topology changes at the limit.
    The extremal capped black hole has two endpoints at $\tilde{\nu}\to 1$ and $\tilde{\gamma}\to 1$ ($A$ in Fig. 4) and $\tilde{\nu}\to\sqrt{2}$ and $\tilde{\gamma}\to 0$ ($B$ in Fig. 4), which have the dimensionless angular momenta and entropy, respectively,
    $(j_{\psi},j_{\phi})\bigr{|}_{A}$ $=\frac{(j_{1}+\sigma^{3}j_{2},j_{2}+\sigma^{3}j_{1})}{(1+\sigma^{2})^{3/2}}\biggr{|}_{A},$ (100)
    $j_{1}\bigr{|}_{A}$ $=1+j_{2}\bigr{|}_{A}=\biggr{\{}\begin{array}[]{cc}1-9^{-1}(2-\sigma^{3})^{2}&(2^{-1/3}<\sigma<1)\\ 1-\sigma^{6}&(-1<\sigma<2^{-1/3})\end{array},$ (103)
    $a_{{\cal H}}\bigr{|}_{A}$ $=2\sqrt{2}\left(\frac{1+\sigma^{3}}{1+\sigma^{2}}\right)^{3/2}\times\quad\biggr{\{}\begin{array}[]{cc}9^{-1}(2-\sigma^{3})\sqrt{2\sigma^{3}-1}&(2^{-1/3}<\sigma<1)\\ \sqrt{\sigma^{6}(1-2\sigma^{3})}&(-1<\sigma<2^{-1/3})\end{array}.$ (106)
    and
    $(j_{\psi},j_{\phi})\bigr{|}_{B}$ $=\frac{(1,\sigma^{3})}{(1+\sigma^{2})^{3/2}},$ (107)
    $a_{{\cal H}}\bigr{|}_{B}$ $=0.$ (108)
    In particular, the endpoint $B$ corresponds to the extremal Cvetič-Youm phase with zero entropy.
  • Nonsupersymetric, extremal capped black hole limit (black curves in Fig. 4) as $\tilde{\nu}\to 1$ and $\tilde{\gamma}\to 1$ with
    $\tilde{\gamma}=1-(\delta^{2}/\mu^{2})\beta,\quad\tilde{\nu}=1+\delta^{2}\beta,\quad(0<\mu<1).$ (109)
    $(\sigma,\tilde{b},\delta,\mu)$ which can be solved for a given set of $(\sigma,\mu)$. Substituting these expressions into Eqs. (84) and (87), and subsequently taking the limit $\beta\to 0$, we derive these two equations for the parameters $(\sigma,\tilde{b},\delta,\mu)$, solvable for a given $(\sigma,\mu)$ configuration. This phase intersects with the previous extremal black hole limit (orange solid curves) at $\mu\to 1$ ($A$ in Fig. 4) and terminates at $\mu\to 0$ ($C$ in Fig. 4). At the latter point, the spacetime corresponds to the extremal phase with zero entropy. For instance, at $\sigma=0.8$, the physical parameters are numerically determined as follows:
    $(j_{\psi},j_{\phi},\phi_{{\mathcal{D}}},q_{{\mathcal{D}}})\bigr{|}_{C}=(0.482...,0.240...,0.780...,-0.0402...).$ (110)
  • Critical curve (red dot-dashed curves in Fig. 4), which is defined by the condition
    $\partial_{\tilde{\nu}}j_{\psi}\partial_{\tilde{\gamma}}j_{\phi}-\partial_{\tilde{\nu}}j_{\phi}\partial_{\tilde{\gamma}}j_{\psi}=0,$ (111)
    where $j_{\psi}$ has a local minimum on a $j_{\phi}={\rm constant}$ surface. This curve forms part of the boundary in the phase diagram only in the $(j_{\psi},j_{\phi})$ plane, as shown in Fig. 4(a). In the $(\phi_{\mathcal{D}},j_{\phi})$ and $(q_{\mathcal{D}},j_{\phi})$ planes, it distinguishes between the large and small bubble branches, as illustrated Fig. 4(b) and (c). This curve ends on the curve $\tilde{\gamma}=0$ ($D$ in Fig. 4) and on the curve of the extremal black hole limit ($E$ in Fig. 4). The respective physical quantities at these end points are numerically determined as
    $(j_{\psi},j_{\phi},\phi_{{\mathcal{D}}},q_{{\mathcal{D}}})\bigr{|}_{D}$ $=(0.461...,0.252...,0.817...,-0.00744...),$ (112)
    $(j_{\psi},j_{\phi},\phi_{{\mathcal{D}}},q_{{\mathcal{D}}})\bigr{|}_{E}$ $=(0.420...,0.187...,0.866...,0).$ (113)
    The $\tilde{\gamma}=0$ ($\gamma=1)$ limit (black dashed curves in Fig. 4). This is a nonextremal and nonsigular phase. This allows the potential possibility to extend the phase diagram beyond this curve, as discussed in Ref. Suzuki:2024vzq for the charge black ring. In such an extended region, $\gamma$ might exceed $1$, leading $(\beta,a,\tilde{b})$ to assume complex values, which we do not discuss here. This curve initiates at $\tilde{\nu}=\sqrt{2}$ at point $B$ in Fig. 4 and concludes at $\tilde{\nu}=1$ at point $C^{\prime}$ in Fig 4(c), where the parameters are given by:
    $(j_{\psi},j_{\phi},\phi_{{\mathcal{D}}},q_{{\mathcal{D}}})\bigr{|}_{C^{\prime}}=(0.482\ldots,0.240\ldots,0.789\ldots,-0.0430\ldots).$ (114)
  • The blue dot-dashed curve in Fig. 4(c) defined by the condition:
    $\partial_{\tilde{\nu}}j_{\phi}\partial_{\tilde{\gamma}}q_{\mathcal{D}}-\partial_{\tilde{\nu}}q_{\mathcal{D}}\partial_{\tilde{\gamma}}j_{\phi}=0,$ (115)
    where $q_{\mathcal{D}}$ has a local maximum on the $j_{\phi}={\rm constant}$ phase (Fig. 5). As in Fig. 4(c), this curve forms a part of the boundary in the $(q_{\mathcal{D}},j_{\phi})$ plane, stretching from the point $C^{\prime}$ to the point where it intersects with the curve of the extremal capped black hole limit, illustrated as $F$ in Fig. 4(c), where
    $(j_{\psi},j_{\phi},\phi_{{\mathcal{D}}},q_{{\mathcal{D}}})\bigr{|}_{F}=(0.481...,0.239...,0.781...,-0.0398...).$ (116)
    Figure 5: $q_{{\mathcal{D}}}$ on the $j_{\phi}=0.2404$ phase near the extremal capped black hole limit (black dot). The local maximum (blue dot) corresponds to the blue dot-dashed curve in Fig. 4(c).
Figure 6: Bubble area of large (blue) and small (red) bubble branches for $\sigma=0.8$ and $j_{\phi}=0.235$.
IV.3 Nonuniqueness of spherical black holes: large and small bubble branches
In our previous works Suzuki:2023nqf ; Suzuki:2024phv , we have demonstrated that the capped black hole with three independent parameters can carry conserved charges–mass, two angular momenta, and electric charge– identical to those of the Cvetič-Youm black hole. This reveals a level of non-uniqueness among spherical black holes, extending beyond the known non-uniqueness attributed to the existence of different topological black objects such as black holes and black rings. Furthermore, the present capped black hole with four independent parameters surprisingly possesses further two branches: one with a large bubble and another with a small bubble, yet both admitting the same conserved charges.
As in Fig. 4, the capped black hole solution displays discrete non-uniqueness relative to the charges $(M,J_{\psi},J_{\phi},Q)$, and supports two distinct solutions: large bubble branch and small bubble branch, differentiated by the relative sizes of their bubble areas. Specifically, the large bubble branch features a greater bubble area compared to its counterpart. This difference is quantified using the dimensionless bubble area defined as:
$a_{\mathcal{D}}:=\frac{A_{\mathcal{D}}}{2\pi r_{M}^{2}},\quad$ (117)
where $A_{\mathcal{D}}$ is defined in Eq. (50). In Fig. 6, we compare the bubble area of the large and small branches for a given $(\sigma,j_{\psi},j_{\phi})$.
Figure 7 compares the entropy of the large, small bubble branches, and the Cvetič-Youm black hole of the same mass-to-charge ratio given by Eq. (88) with $\sigma=0.8$ having same angular momenta $(j_{\psi},j_{\phi}=0.235)$, where we label the values of $j_{\psi}$ at the cross section of the extremal capped black hole limit, extremal black hole limit, and the critical curve as $j_{\psi,{\rm ext}}^{\rm cap}$, $j_{\psi,{\rm ext}}^{\rm CY}$, and $j_{\psi,c}$, respectively. We observe that three distinct solutions are present for $j_{\psi,c}<j_{\psi}<j_{\psi,{\rm ext}}^{\rm CY}$, whereas only the large bubble branch exists for $j_{\psi,{\rm ext}}^{\rm CY}<j_{\psi}<j_{\psi,{\rm ext}}^{\rm cap}$. The area where three solutions are possible corresponds to the purple region in Fig. 4(a). In Fig. 4(b) and (c), the small bubble branch is depicted in red, and the corresponding large bubble branch is depicted in blue. These regions are separated from the region admitting only the large bubble branch (gray region in Fig. 4) by the curve $j_{\psi}=j_{\psi}^{\rm BH}(\sigma,j_{\phi})$ (orange dashed curves in Fig. 4(b) and (c)) where $j_{\psi}^{\rm BH}$ is defined by Eq. (95).
Moreover, the value $j_{\psi,\times}$ can be identified as the phase where the entropies of the Cvetič-Youm black hole and the large bubble branch are equal. For $j_{\psi,c}<j_{\psi}<j_{\psi,\times}$ , the Cvetič-Youm black hole is thermodynamically favored due to its higher entropy, while for $j_{\psi,\times}<j_{\psi}<j_{\psi,{\rm ext}}^{\rm CY}$, the large bubble branch has the greater entropy and is thus preferred. The small bubble branch occasionally exhibits higher entropy than the large bubble branch near the critical curve, but it remains thermodynamically irrelevant as it has lower entropy compared to the Cvetič-Youm black hole.
Figure 7: Entropy of large and small bubble branches for $\sigma=0.8$ and $j_{\phi}=0.235$ displayed by the blue and red curves, respectively. The entropy of the Cvetič-Youm black hole with the same charges is also shown by the gray curve.
IV.4 Relation with three parameter solution Suzuki:2023nqf ; Suzuki:2024phv
Finally, we discuss the connection between the new four-parameter solution and the previously three-parameter capped black hole solution, as detailed in Refs.Suzuki:2023nqf ; Suzuki:2024phv . This three-parameter solution is derived from the four-parameter model by simply setting $\beta=0$. As shown in Fig. 8, the $\beta=0$ phase is situated near the $\tilde{\gamma}=0$ curve. It begins at the singular extremal Cvetič-Youm phase (marked as $B$ in Fig. 8) with $\tilde{\nu}=\sqrt{2}$ and ends at the zero entropy extremal capped black hole phase (marked as $C^{\prime}$ in Fig. 8) with $\tilde{\nu}=1$. Notably, within the three-parameter family, the $\beta=0$ phase does not exhibit overlaps in the $(j_{\psi},j_{\phi})$-plane, thus not displaying the non-uniqueness associated with the four-parameter family.
Figure 8: The phase of the three parameter capped black hole solution (black curve) embedded in the four parameter phase space for $\sigma=0.8$ presented in Fig. 4.
V Summary
In this paper, we have constructed the new non-BPS solution of the capped black hole in the bosonic sector of five-dimensional minimal supergravity, which describes an asymptotically flat, stationary, bi-axisymmetric, charged spherical black hole with the DOC of the topology of $[\mathbb{R}^{4}\#\mathbb{CP}^{2}]\setminus{\mathbb{B}}^{4}$ on a timeslice. Using the Harrison transformation, we have previously constructed the three-parameter family solution of the capped black hole Suzuki:2023nqf ; Suzuki:2024phv . By combining the Ehlers and Harrison transformations, we could have generalized it to the four-parameter family solution, allowing all conserved charges to be independent. We have found that the new solution displays discrete non-uniqueness with respect to the conserved charges, allowing two distinct branches for the same mass, angular momenta, and electric charge. These branches, called the large bubble and small bubble branches, differ in the size of the bubble adjacent to the horizon. We have also found that the Cvetič-Youm black hole exists for the conserved charges with which two capped black hole branches exist. Whereas, with the angular momenta beyond the extremal limit of the Cvetič-Youm black hole, there exists only the large bubble branch. Our analysis have shown that, among the three, the most thermodynamically preferred solution is the large bubble branch if the bubble size is sufficiently large, but otherwise the Cvetič-Youm black hole. While the small bubble branch always has smaller entropy than the Cvetič-Youm black hole, it can have larger entropy than the large bubble branch, at least for $\sigma=0.8$ near the critical curve.
The phase diagram we have obtained may not complete since the parameter region is partially bounded by nonextremal, nonsingular phases, indicating the possible extension of the parameter region as in Ref. Suzuki:2024vzq . In the extended region, some parameters would take complex values, even if the metric and gauge potential remain real. It may also be an interesting issue to construct non-BPS black hole solutions with the DOC of other non-trivial topology, $[{\mathbb{R}}^{4}\#n{\mathbb{C}}{\mathbb{P}}^{2}\#m(S^{2}\times S^{2})]\setminus{\mathbb{B}}^{4}$ with $(n,m)\not=(1,0)$ on a timeslice.
Acknowledgement
R.S. was supported by JSPS KAKENHI Grant Number JP24K07028. S.T. was supported by JSPS KAKENHI Grant Number 21K03560.
Appendix A Coefficients $f_{i}$ and $g_{i}$
The explicit forms of the coefficients $f_{i}$ and $g_{i}$ are shown below.
A.1 $f_{1},\dots,f_{19}$
Every $f_{i}$ is a quadratic function of $\beta$. $f_{i}$ for $16\leq i\leq 19$ appears only in the expressions of $g_{i}$.
$f_{1}$ $=1+2(a-b)\beta+\frac{\beta^{2}}{\gamma^{2}-\nu^{2}}\biggr{[}(a-b)^{2}\gamma^{2}-(1-a)^{2}\gamma(1+2b+(3+2b)\nu)$
$\quad\quad+\nu\left((1-a)^{2}(3+2b)+\left(1-2a+2\left(1-a+a^{2}\right)b-b^{2}\right)\nu\right)\biggr{]},$ (118)
$f_{2}$ $=1+2(1-ab)\beta+\frac{\beta^{2}}{\gamma^{2}-\nu^{2}}\biggr{[}\left((2-a)a-2\left(1-a+a^{2}\right)b-(1-2a)b^{2}\right)\gamma^{2}$
$\quad\quad+(1-a)^{2}\gamma(1+2b+(3+2b)\nu)-\nu\left((1-a)^{2}(3+2b)+(1-b)(1+(1-2a)b)\nu\right)\biggr{]},$ (119)
$f_{3}$ $=1+2(a-b)\beta+\frac{\beta^{2}}{\gamma^{2}-\nu^{2}}\biggr{[}(a-b)^{2}\gamma^{2}-2(1-a)^{2}(1+a)+(1-a)^{2}(1-2a+2b)\nu$
$\quad\quad+\left(1-2a+2\left(1-a+a^{2}\right)b-b^{2}\right)\nu^{2}+(1-a)^{2}\gamma(1+2a-2b-(1-2a+2b)\nu)\biggr{]},$ (120)
$f_{4}$ $=1-\frac{2\beta}{\gamma-\nu}\left(a^{2}-1-a(a-b)\gamma+\nu(1-ab)\right)+\frac{\beta^{2}}{\gamma^{2}-\nu^{2}}\biggr{[}2(1-a)^{2}(1+a)-(1-2a)(a-b)^{2}\gamma^{2}$
$\quad\quad-(1-a)(1-5a+4b)\nu-\left(1-2ab-b^{2}+2ab^{2}\right)\nu^{2}-(1-a)\gamma\left(1-a-4a^{2}+4ab-(1-5a+4b)\nu\right)\biggr{]},$ (121)
$f_{5}$ $=1-\frac{2\beta}{c_{2}}\left(2a(-1+\gamma)\nu+ab^{2}(\gamma-\nu)(1+\nu)+b\left(-1+a^{2}-a^{2}\gamma-\left(2+a^{2}\right)(-1+\gamma)\nu+\nu^{2}\right)\right)$
$\quad+\frac{\beta^{2}}{(\gamma^{2}-\nu^{2})c_{2}}\biggr{[}b\gamma\left(2(1-a)^{2}(1+a)+(1-a)\left(-1+a+4a^{2}-4ab\right)\gamma+(-1+2a)(-a+b)^{2}\gamma^{2}\right)$
$\quad+\nu\biggr{\{}\left(2\left(-2+a+6a^{2}-4a^{3}\right)+\left(4-20a+17a^{2}-6a^{3}\right)b+4\left(2-3a+2a^{2}\right)b^{2}+(1-2a)b^{3}\right)\gamma^{2}$
$\quad\quad+(-a+b)^{2}(-4+2a-b+2ab)\gamma^{3}-2(1-a)^{2}(1+a)(2+b)+2(1-a)^{2}\left(4+5a+(-1+3a)b-2b^{2}\right)\gamma\biggr{\}}$
$\quad+\nu^{2}\biggr{\{}(1-a)\left(2\left(2-3a+a^{2}\right)+\left(3-9a+2a^{2}\right)b+4b^{2}\right)$
$\quad\quad+\left(4(a-2)(1-a)^{2}+\left(-10+26a-21a^{2}+4a^{3}\right)b-2\left(4-7a+2a^{2}\right)b^{2}+(1-2a)b^{3}\right)\gamma$
$\quad\quad+\left(2(2-a)(1-a)^{2}+\left(5-14a+10a^{2}-2a^{3}\right)b+\left(4-6a+4a^{2}\right)b^{2}+(1-2a)b^{3}\right)\gamma^{2}\biggr{\}}$
$\quad-\left(2a-\left(6-6a+5a^{2}\right)b+4ab^{2}+(1-2a)b^{3}\right)(1-\gamma)\nu^{3}+(1-b)b(1+(1-2a)b)\nu^{4}\biggr{]},$ (122)
$f_{6}$ $=1-\frac{2\beta}{(\gamma-\nu)(1+\nu)c_{1}}\biggr{[}a^{2}-1+\left(2-2a+a^{2}+b-2ab\right)\nu-(1-b)(1-2a+b)\nu^{2}-(1-b)b\nu^{3}$
$\quad\quad+(a-b)^{2}\gamma^{2}(1+\nu)+\gamma\left(1-2a^{2}-b+2ab-2\left(1-a+a^{2}-2ab+b^{2}\right)\nu+(1-b)(1-2a+2b)\nu^{2}\right)\biggr{]}$
$\quad+\frac{\beta^{2}}{(\gamma^{2}-\nu^{2})(1+\nu)c_{1}}\biggr{[}2(1-a)^{2}(1+a)-(1-a)\left(2-5a-a^{2}+4b\right)\nu$
$\quad\quad+\left(-2-2a+6a^{2}-3a^{3}+\left(3-4a+3a^{2}\right)b+(2-3a)b^{2}\right)\nu^{2}+\left(2-3a+3a^{2}b-3ab^{2}+b^{3}\right)\nu^{3}+(1-b)^{2}b\nu^{4}$
$\quad\quad+(a-b)\gamma^{2}\left(4a^{2}-1-2a+(2-3a)b+\left((2-a)^{2}-2ab+b^{2}\right)\nu-\left(3(1-a)^{2}+(2-a)b-b^{2}\right)\nu^{2}\right)$
$\quad\quad-(a-b)^{3}\gamma^{3}(1+\nu)+(1-a)\gamma\left(a-2+5a^{2}-4ab+\left(2-9a-a^{2}+8b\right)\nu\right)$
$\quad\quad+\left(2+5a-12a^{2}+6a^{3}-\left(8-12a+7a^{2}\right)b+3ab^{2}-b^{3}\right)\gamma\nu^{2}-\left(2-3a+3a^{2}b-3ab^{2}+b^{3}\right)\gamma\nu^{3}\biggr{]},$ (123)
$f_{7}$ $=1-\frac{2\beta}{b(\gamma-\nu)(1+\nu)}\biggr{[}(b-1)b\gamma+\left(2-2a+b-b^{2}\right)(1-\gamma)\nu+(1-b)b\nu^{2}\biggr{]}$
$+\frac{\beta^{2}}{b(\gamma^{2}-\nu^{2})(1+\nu)}\biggr{[}b\gamma\left((1-a)^{2}+\left((2-a)a-2b+b^{2}\right)\gamma\right)-(1-b)^{2}b\nu^{3}$
$\quad+\left(-(1-a)^{2}(4+3b)+4(1-a)(1-2a+(2-a)b)\gamma+\left(4(1-a)a-\left(4-6a+a^{2}\right)b-2b^{2}+b^{3}\right)\gamma^{2}\right)\nu$
$\quad+\left(4(-1+a)+(2-3a)ab+2b^{2}-b^{3}+(1-a)(4-b-3ab)\gamma\right)\nu^{2}\biggr{]},$ (124)
$f_{8}$ $=1+\frac{2\beta}{(a-b)(\gamma-\nu)}\biggr{[}1-a^{2}+(a-b)^{2}\gamma-\left(1-2ab+b^{2}\right)\nu\biggr{]}$
$-\frac{\beta^{2}}{(a-b)\left(\gamma^{2}-\nu^{2}\right)}\biggr{[}2(-1+a+a^{2}-a^{3})-a\gamma(1+2a-3a^{2})-(a^{3}-b^{3})\gamma^{2}-b^{2}\gamma\left(2+2a^{2}+a(-4+3\gamma)\right)$
$\quad+b\left(-2(1-a)^{2}(1+a)+\left(1+4a-7a^{2}+2a^{3}\right)\gamma+3a^{2}\gamma^{2}\right)$
$\quad-\nu(1-a)(\gamma-1)\left((a-5)a+\left(5-3a+2a^{2}\right)b+2(1-a)b^{2}\right)$
$\quad-\left(-2+a+\left(-1+6a-2a^{2}\right)b+\left(-2+a-2a^{2}\right)b^{2}+b^{3}\right)\nu^{2}\biggr{]},$ (125)
$f_{9}$ $=1+2(a-b)\beta+\frac{\beta^{2}}{\left(\gamma^{2}-\nu^{2}\right)c_{1}}\biggr{[}-2(1-a)^{2}(1+a)-(a-b)^{3}\gamma^{3}+(1-a)^{2}(4+a+2b)\nu-(1-b)b(1-2a+b)\nu^{3}$
$-\left(2-5a+4a^{2}+\left(1-2a-a^{2}\right)b+ab^{2}\right)\nu^{2}-(a-b)\gamma^{2}\left(1-a(2-b)-\left((1-a)^{2}-ab+b^{2}\right)\nu\right)$
$-\gamma\left((1-a)^{2}(-2-3a+2b)+2\left(2-3a+a^{3}\right)\nu+\left(-2+5a-4a^{2}+\left(-2+4a+a^{2}\right)b-3ab^{2}+b^{3}\right)\nu^{2}\right)\biggr{]},$ (126)
$f_{10}$ $=\frac{4(1-a)(1-\gamma)d_{1}}{\gamma^{2}-\nu^{2}}\beta(1+(a-b)\beta),$ (127)
$f_{11}$ $=1-\frac{2\beta(-1+a-a\gamma+b(-1+a+\gamma)+(-1+ab)\nu)}{\gamma+\nu}$
$+\frac{\beta^{2}}{\gamma^{2}-\nu^{2}}\biggr{[}(a-b)^{2}\gamma^{2}+\nu\left((1-a)\left(a-3+2b^{2}\right)-(1-b)(1+b-2ab)\nu\right)+(1-a)\gamma\left(1+a-2b^{2}+\left(3-a-2b^{2}\right)\nu\right)\biggr{]},$ (128)
$f_{12}$ $=f_{1}+\frac{c_{3}\left(f_{1}-f_{3}\right)}{(a+1)b(1-\gamma)(\nu+1)},$ (129)
$f_{13}$ $=f_{7}+\frac{\left(b(\gamma-\nu^{2})-(2+b)(1-\gamma)\nu\right)f_{10}}{2b(1+\nu)d_{1}},$ (130)
$f_{14}$ $=1-\frac{2\beta}{d_{2}}\biggr{[}b\left(-1+a^{2}+\gamma-2a^{2}\gamma-(1-2a)b\gamma+(a-b)^{2}\gamma^{2}\right)$
$\quad+\left(-2+\left(2+a^{2}\right)b+(1-2a)b^{2}-2\left(-1+b+a^{2}b-2ab^{2}+b^{3}\right)\gamma+b(-a+b)^{2}\gamma^{2}\right)\nu$
$\quad+\left(2-b-2ab^{2}+b^{3}+\left(-2+b+(1+2a)b^{2}-2b^{3}\right)\gamma\right)\nu^{2}+(-1+b)b^{2}\nu^{3}\biggr{]}$
$+\frac{\beta^{2}}{\left(\gamma^{2}-\nu^{2}\right)d_{2}}\biggr{[}b\gamma\left(2(1-a)^{2}(1+a)+(1-a)\left(-2+a+5a^{2}-4ab\right)\gamma\right.$
$\quad\quad\left.+\left(a\left(-1-2a+4a^{2}\right)+\left(1+4a-7a^{2}\right)b+(-2+3a)b^{2}\right)\gamma^{2}-(a-b)^{3}\gamma^{3}\right)$
$\quad+\left(-2(1-a)^{2}(1+a)(2+b)+2(1-a)^{2}\left(5+6a+4ab-2b^{2}\right)\gamma\right.$
$\quad\quad\left.+\left(-6+4a+16a^{2}-12a^{3}+\left(2-22a+26a^{2}-11a^{3}\right)b+\left(9-16a+11a^{2}\right)b^{2}+(2-3a)b^{3}\right)\gamma^{2}\right.$
$\quad\quad\left.-2\left((3-2a)a^{2}-3a\left(2-2a+a^{2}\right)b+\left(3-4a+5a^{2}\right)b^{2}-3ab^{3}+b^{4}\right)\gamma^{3}+b(-a+b)^{3}\gamma^{4}\right)\nu$
$\quad+\left((1-a)\left(2-6a+4a^{2}+\left(2-9a+3a^{2}\right)b+4b^{2}\right)\right.$
$\quad\quad+\left(6(1-a)^{2}(-1+2a)+\left(-12+37a-34a^{2}+8a^{3}\right)b+\left(-13+20a-5a^{2}\right)b^{2}+(2-3a)b^{3}\right)\gamma$
$\quad\quad+\left(4-24a+30a^{2}-12a^{3}+\left(18-39a+32a^{2}-7a^{3}\right)b+2\left(5-10a+4a^{2}\right)b^{2}\right)\gamma^{2}$
$\quad\quad\left.+\left(2a\left(4-5a+2a^{2}\right)+\left(-8+13a-10a^{2}+2a^{3}\right)b+\left(-3+4a-3a^{2}\right)b^{2}+(2+3a)b^{3}-2b^{4}\right)\gamma^{3}\right)\nu^{2}$
$\quad+\left(2(a-2)a-\left(-10+14a-10a^{2}+a^{3}\right)b+\left(3-8a+a^{2}\right)b^{2}+(-2+3a)b^{3}\right.$
$\quad\quad\left.+2\left(-1+4a-2a^{2}+\left(-8+14a-10a^{2}+a^{3}\right)b-\left(3-8a+a^{2}\right)b^{2}-3ab^{3}+b^{4}\right)\gamma\right.$
$\quad\quad\left.+\left(2(-1+a)^{2}-\left(-6+14a-10a^{2}+a^{3}\right)b+\left(5-8a+a^{2}\right)b^{2}+(-2+3a)b^{3}\right)\gamma^{2}\right)\nu^{3}-(1-b)^{2}b^{2}\nu^{5}$
$\quad+\left(2-(2+a)b+\left(2-4a+a^{2}\right)b^{2}+3ab^{3}-b^{4}+\left(-2+(2+a)b-\left(1-4a+a^{2}\right)b^{2}-(2+3a)b^{3}+2b^{4}\right)\gamma\right)\nu^{4}\biggr{]},$ (131)
$f_{15}$ $=1-\frac{2\beta}{(\gamma-\nu)\left(c_{3}-(1-ab)(1-\gamma)(1+\nu)\right)}\biggr{[}1-a^{2}-\left(2+a^{2}+a\left(-2-2b+b^{2}\right)\right)\nu$
$+\left(1-2(1-a)b-(1+a)b^{2}+b^{3}\right)\nu^{2}-(1-b)b^{2}\nu^{3}-(a-b)\gamma^{2}\left(a-b+b^{2}+\left(-2+a-b+b^{2}\right)\nu\right)$
$+\gamma\left(-1+2a^{2}-2ab+ab^{2}+2\left((1-a)^{2}+b-2ab+(1+a)b^{2}-b^{3}\right)\nu+\left(-1+2(1-a)b+(2+a)b^{2}-2b^{3}\right)\nu^{2}\right)\biggr{]}$
$+\frac{\beta^{2}}{(\gamma^{2}-\nu^{2})\left(c_{3}-(1-ab)(1-\gamma)(1+\nu)\right)}\biggr{[}-2(1-a)^{2}(1+a)(1+b)$
$\quad-(1-a)\left(1+5a-2a^{2}+\left(-4+3a-3a^{2}\right)b-2(1-a)b^{2}\right)\nu$
$\quad+\left(4-6a+3a^{2}-\left(-2+8a-4a^{2}+a^{3}\right)b+\left(2-2a+3a^{2}\right)b^{2}-ab^{3}\right)\nu^{2}$
$\quad+\left(-1+(4-3a)b+\left(-1+2a+a^{2}\right)b^{2}-(2+a)b^{3}+b^{4}\right)\nu^{3}+(-1+b)^{2}b^{2}\nu^{4}$
$\quad+(a-b)^{2}\gamma^{3}\left(-1+2a+(-2+a)b+b^{2}+\left(-3+2a+(-2+a)b+b^{2}\right)\nu\right)$
$\quad+\gamma^{2}\left(-1+2a+4a^{2}-6a^{3}-\left(2+5a-14a^{2}+4a^{3}\right)b+\left(4-10a+3a^{2}\right)b^{2}+ab^{3}\right.$
$\quad\quad\left.+\left(-2-4a+13a^{2}-6a^{3}+\left(8-22a+18a^{2}-5a^{3}\right)b+\left(7-14a+5a^{2}\right)b^{2}+(2+a)b^{3}-b^{4}\right)\nu\right.$
$\quad\quad\left.+\left(3(-1+a)^{2}-(-2+a)(-1+a)^{2}b+\left(1-4a+2a^{2}\right)b^{2}+2b^{3}-b^{4}\right)\nu^{2}\right)$
$\quad+\gamma\left((-1+a)\left(-3+a+6a^{2}+\left(-4-5a+5a^{2}\right)b-2(-1+a)b^{2}\right)\right.$
$\quad\quad\left.+(-1+a)\left(-3-11a+6a^{2}+\left(12-11a+7a^{2}\right)b-6(-1+a)b^{2}\right)\nu\right.$
$\quad\quad\left.-\left(7-12a+6a^{2}+\left(4-13a+8a^{2}-2a^{3}\right)b+\left(5-6a+5a^{2}\right)b^{2}-(2+a)b^{3}+b^{4}\right)\nu^{2}\right.$
$\quad\quad\left.+\left(1+(-4+3a)b-\left(-1+2a+a^{2}\right)b^{2}+(2+a)b^{3}-b^{4}\right)\nu^{3}\right)\biggr{]},$ (132)
$f_{16}$ $=1-\frac{2\beta\left(-1+a^{2}+(a-b)^{2}\gamma^{2}+(1-2a+b)\nu+(1-b)b\nu^{2}-\gamma\left(-1+2a^{2}+b-2ab+(1-2a+b)\nu\right)\right)}{(\gamma+\nu)c_{1}}$
$+\frac{\beta^{2}}{\left(\gamma^{2}-\nu^{2}\right)c_{1}}\biggr{[}2(1-a)^{2}(1+a)-(a-b)^{3}\gamma^{3}-(1-a)\left(4+a^{2}-a(1+4b)\right)\nu$
$\quad+\left(2-3a-\left(1-4a+a^{2}\right)b+(-2+a)b^{2}\right)\nu^{2}+(1-b)^{2}b\nu^{3}$
$\quad+(a-b)\gamma^{2}\left(-1-2a+4a^{2}+(2-3a)b+\left((-1+a)^{2}+(2-3a)b+b^{2}\right)\nu\right)$
$\quad+\gamma\left((1-a)\left(-2+a+5a^{2}-4ab\right)+2(1-a)\left(2-a+a^{2}-4ab+2b^{2}\right)\nu\right)$
$\quad-\left(2-3a+(4-a)ab-(4-a)b^{2}+b^{3}\right)\gamma\nu^{2}\biggr{]},$ (133)
$f_{17}$ $=1+\frac{2\beta(1-a+(a-b)\gamma+(1-b)\nu)}{\gamma+\nu}$
$+\frac{\beta^{2}\left((a-b)^{2}\gamma^{2}+\nu\left((1-a)(1-3a+2b)-(1-b)^{2}\nu\right)+(1-a)\gamma(1+a-2b-\nu+3a\nu-2b\nu)\right)}{\gamma^{2}-\nu^{2}},$ (134)
$f_{18}$ $=1+\frac{2\beta\left(1-a^{2}+a(a-b)\gamma+(2-a)(a-b)(1-\gamma)\nu-(1-ab)\nu^{2}\right)}{(1-\nu)(\gamma+\nu)}$
$+\frac{\beta^{2}}{(\gamma^{2}-\nu^{2})(1-\nu)}\biggr{[}2(1-a)^{2}(1+a)-(1-a)\left(3+2a^{2}-a(1+4b)\right)\nu+\left((a-2)a+2(3-2a)ab-(3-2a)b^{2}\right)\nu^{2}$
$\quad+(1-b)(1+(1-2a)b)\nu^{3}+(a-b)^{2}\gamma^{2}(-1+2a+(-3+2a)\nu)$
$\quad+(1-a)\gamma\left(-1+a+4a^{2}-4ab+\left(2-2a+4a^{2}-8ab+4b^{2}\right)\nu+\left(-1+a-4ab+4b^{2}\right)\nu^{2}\right)\biggr{]},$ (135)
$f_{19}$ $=1+\frac{2\beta(1-a+(a-b)\gamma-(1-b)\nu)}{\gamma-\nu}+\frac{\beta^{2}\left(1-a^{2}-2b(1-a(1-\gamma))+a^{2}\gamma+b^{2}\gamma-(1-b)^{2}\nu\right)}{\gamma-\nu}.$ (136)
A.2 $g_{1},\dots,g_{13}$ and $\tilde{g}_{1}$
For the metric functions, we have
$g_{1}$ $=f_{13}f_{14}-\frac{c_{1}c_{3}\left(f_{13}f_{14}-f_{16}f_{19}\right)}{2(a-b)(1-\gamma)^{2}\nu}+\frac{(1-\nu)c_{3}f_{3}f_{10}}{2(a-b)(1-\gamma)(1+\nu)d_{1}},$ (137)
$g_{2}$ $=\Delta+\frac{c_{3}\left(\Delta-f_{3}f_{4}\right)}{d_{3}},$ (138)
$g_{3}$ $=f_{14}f_{18}-\frac{d_{1}\left(f_{14}f_{18}-f_{2}f_{17}\right)}{(1-\gamma)(1+\nu)\left(ac_{1}-1+\nu\right)}-\frac{b\nu(\gamma-\nu)f_{10}f_{12}}{(1-\gamma)(1-\nu^{2})\left(ac_{1}-1+\nu\right)},$ (139)
$g_{4}$ $=\Delta+\frac{1}{d_{4}}\biggr{[}-(1-\gamma)^{2}(1-\nu)^{2}\nu d_{1}\left(\Delta-f_{1}^{2}\right)-\frac{8\nu d_{1}\left(d_{1}-(1+a)(1-\gamma)(1+\nu)c_{1}\right)\left(\Delta-f_{1}f_{2}\right)}{1+a}$
$+2(1-\gamma)(1-\nu)\nu\left(-2-3\nu+3\nu^{2}\right)d_{1}\left(\Delta-f_{2}^{2}\right)-\frac{1}{2}(1-\nu)(3-\nu)d_{2}^{2}\left(\Delta-f_{14}^{2}\right)$
$-\frac{2(\gamma-\nu)(1-\nu)^{3}c_{3}\left(c_{3}+b(1+a)(1-\gamma)(1+\nu)\right)\left(\Delta-f_{3}f_{4}\right)}{1+a}+2(1-\gamma)(1-\nu)^{2}\nu c_{2}^{2}\left(\Delta-f_{5}^{2}\right)$
$+\frac{1}{2}(\gamma-\nu)(1+\nu)\left(3-2\nu+\nu^{2}\right)c_{1}^{2}c_{3}\left(\Delta-f_{6}^{2}\right)+(1-\nu)\nu^{2}\left(c_{3}-2(1-\gamma)\nu\right)d_{1}\left(\Delta-f_{7}^{2}\right)$
$+2(a-b)^{2}(1-\gamma)^{3}(1-\nu)(\gamma-\nu)\nu^{3}\left(\Delta-f_{8}^{2}\right)+\frac{(\gamma-\nu)(-1+\nu)^{2}\nu^{2}(\gamma+\nu)c_{3}f_{10}^{2}}{(1+\nu)d_{1}}\biggr{]},$ (140)
$g_{5}$ $=\Delta-\frac{d_{1}\left(\Delta-f_{1}f_{2}\right)}{ad_{1}+\left(1-a^{2}\right)(1-\gamma)(1+\nu)c_{1}},$ (141)
$g_{6}$ $=f_{8}^{2}+\frac{2(1-\gamma)^{2}(1-\nu)^{2}\nu d_{1}\left(f_{1}^{2}-f_{8}^{2}\right)}{(1+\nu)d_{5}}+\frac{(1-\gamma)^{2}(1-\nu)^{4}c_{3}\left(f_{3}^{2}-f_{8}^{2}\right)}{(1+\nu)d_{5}},$ (142)
$g_{7}$ $=g_{8}+\frac{c_{3}f_{10}^{2}(\nu-1)^{2}\nu\left(\gamma^{2}-\nu^{2}\right)}{d_{1}d_{6}(\nu+1)},$ (143)
$g_{8}$ $=f_{14}^{2}-\frac{2(1-\gamma)(1-\nu)^{2}\nu d_{1}\left(f_{14}^{2}-f_{2}^{2}\right)}{d_{6}}+\frac{(\gamma-\nu)(1-\nu)^{2}(1+\nu)c_{1}^{2}c_{3}\left(f_{14}^{2}-f_{6}^{2}\right)}{2d_{6}\nu},$ (144)
$g_{9}$ $=f_{8}f_{18}+\frac{(2+\nu)f_{1}f_{10}}{(a-b)(1-\gamma)(1-\nu^{2})},$ (145)
$g_{10}$ $=f_{14}f_{18}-\frac{bf_{10}f_{12}(\nu+1)(\gamma-\nu)}{2d_{2}(\nu-1)},$ (146)
$g_{11}$ $=f_{16}f_{15}-\frac{(1-\gamma)\left(4+\nu-\nu^{2}\right)f_{1}f_{10}}{(1+\nu)\left(2(a-b)(1-\gamma)^{2}(1-\nu)\nu-b(1+\nu)d_{1}\right)}-\frac{(a-b)(1-\gamma)^{2}(1-\nu)^{2}\left(f_{18}f_{8}-f_{16}f_{15}\right)}{2(a-b)(1-\gamma)^{2}(1-\nu)\nu-b(1+\nu)d_{1}}.$ (147)
Additionally, we also define
$\tilde{g}_{1}$ $=g_{1}+\frac{\nu c_{3}f_{3}f_{10}}{(1-\gamma)d_{1}(\nu+1)(a-b)}$
$=f_{13}f_{14}+\frac{c_{3}f_{3}f_{10}}{2(a-b)(1-\gamma)d_{1}}-\frac{c_{1}c_{3}\left(f_{13}f_{14}-f_{16}f_{19}\right)}{2(a-b)(1-\gamma)^{2}\nu},$ (148)
$g_{12}$ $=f_{5}f_{13}+\frac{c_{3}(f_{4}f_{17}-f_{5}f_{13})}{2\nu(1-\gamma)(1-ab)},$ (149)
$g_{13}$ $=f_{8}f_{18}-\frac{f_{1}f_{10}}{2(1-\gamma)(1-\nu)(b-a)}.$ (150)